### v3

# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

Abstract.
The healthcare landscape is evolving, with patients seeking reliable information about their health conditions and available treatment options. Despite the abundance of information sources, the digital age overwhelms individuals with excess, often inaccurate information. Patients primarily trust medical professionals, highlighting the need for expert-endorsed health information. However, increased patient loads on experts has led to reduced communication time, impacting information sharing. To address this gap, we develop CataractBot, an experts-in-the-loop chatbot powered by LLMs, in collaboration with an eye hospital in India. CataractBot answers cataract surgery related questions instantly by querying a curated knowledge base, and provides expert-verified responses asynchronously. It has multimodal and multilingual capabilities. In an in-the-wild deployment study with 55 participants, CataractBot proved valuable, providing anytime accessibility, saving time, accommodating diverse literacy levels, alleviating power differences, and adding a privacy layer between patients and doctors. Users reported that their trust in the system was established through expert verification. Broadly, our results could inform future work on designing expert-mediated LLM bots.
1. Introduction
The evolving landscape of healthcare witnesses a significant shift with patients assuming more proactive roles in their care journeys (Klasnja et al., 2010a; Kinnersley et al., 2007), resulting in an increased demand for information. Patients and their caregivers seek accessible, comprehensive, and reliable information about their symptoms, diagnoses, treatment options, potential risks, and preventive measures (Christalle et al., 2019). To satisfy these information needs, patients actively explore diverse sources, including online resources, friends and family, support groups, and direct communication with healthcare professionals (Skeels et al., 2010). In particular, for major treatments like surgery or radiotherapy, satisfying the demand for information becomes crucial (Christalle et al., 2019), with questions spanning pre-, during-, and post-treatment phases. Studies highlight the anxiety patients and caregivers experience regarding such treatments (Klasnja et al., 2010b), emphasizing the correlation between anxiety and negative clinical outcomes (Byrne and Heyman, 1997; Dahlén and Janson, 2002). Similarly, access to information has been found to significantly reduce anxiety (Wilcox et al., 2010) and improve patient satisfaction and clinical outcomes (Kurtz, 2002; Wilcox et al., 2010), thus underlining its pivotal role.
Despite the abundance of available information sources, the digital age often overwhelms individuals with an excess of information, much of which is inaccurate or unreliable (Clarke et al., 2016a; Kinnersley et al., 2007). This poses a challenge in discerning trustworthy sources from misleading ones. Patients also frequently encounter difficulties finding information online at an appropriate level, ranging from oversimplified to overly technical (Xiao et al., 2023). As a result, patients tend to rely on the experiences of friends and family members who have undergone similar treatments, although identifying them can be challenging (Civan-Hartzler et al., 2010). Therefore, patients primarily trust their doctor and the hospital staff responsible for their treatment, turning to them for any medical or logistical queries (Attfield et al., 2006; Clarke et al., 2016a). Studies examining patients’ information needs reveal a strong desire for doctor-endorsed health information (Tang and Newcomb, 1998).
However, the escalating pressure on doctors to accommodate more patients has led to reduced time per patient, affecting communication and information sharing. In developing nations, a lower doctor-to-patient ratio hinders personalized attention and comprehensive guidance (Wang et al., 2021b), while in developed countries, managing Electronic Health Records (EHR) often diminishes direct patient communication quality and available time (Cajander and Grünloh, 2019a). Previous studies (Tongue et al., 2005; Kinnersley et al., 2007) have found that doctors often underestimate patients’ information needs and overestimate the amount of information they provide. However, providing all potentially relevant information risks overwhelming patients with cognitive overload (Christalle et al., 2019). Surveys consistently highlight patients’ desire for improved communication with their healthcare providers (Duffy et al., 2004). This information exchange extends beyond doctor-patient interactions to encompass relationships between various medical professionals and patients, such as patient-nurse interactions (Lotfi et al., 2019; Wang et al., 2020). Moreover, enhanced access to information not only benefits patients but also has the potential to reduce unnecessary visits, subsequently alleviating the burden on doctors and medical staff. To summarize, as Tang and Lansky (2005) correctly stated: “Patients have little access to information and knowledge that can help them participate in, let alone guide, their own care… A simple, non-urgent exchange of questions and answers is often all that is required.”.
To address patients’ information needs, particularly during surgical treatments, we propose a doctor-integrated chatbot solution designed to provide reliable and comprehensive real-time responses. Leveraging the capabilities of large-language models (LLMs) and the widespread use of instant messaging services such as WhatsApp, this chatbot serves as an accessible, 24/7 resource capable of understanding intricate human queries and providing accurate information. In collaboration with a tertiary eye hospital in Bangalore, India, we exemplified our experts-in-the-loop chatbot approach by developing CataractBot. This WhatsApp-based chatbot, powered by LLMs, answers queries related to cataract surgery from patients and their attendants. Recognizing the limitations of generic LLMs in the medical domain, we utilized Retrieval-Augmented Generation (RAG) over a Knowledge Base (KB) curated by doctors and hospital staff to provide hospital-specific and culturally sensitive responses. CataractBot has several features: multimodal support (accepting both speech and text inputs), multilingual capabilities (supporting five languages, including English, Hindi, and Kannada), a verification system where responses are vetted by experts (ophthalmologists for medical queries and patient coordinators for logistical ones), and a KB update framework based on expert-provided edits to minimize future expert intervention. CataractBot integrates both synchronous and asynchronous messaging by providing an LLM-generated answer instantly, and subsequently providing verification at the experts’ convenience.
Our work aims to answer the following research questions: (RQ1) What is the role of an LLM-powered WhatsApp-based chatbot in meeting the information needs of patients (and attendants) undergoing surgery? (RQ2) How do the different features of CataractBot, including experts-in-the-loop and multimodality, contribute to its usage? (RQ3) How can CataractBot be integrated into the current doctor-patient workflow? To answer these questions, we conducted an in-the-wild deployment study involving 49 information seekers111We use the terms ‘participants’ or ‘information seekers’ to refer to patients and their attendants, and ‘experts’ to refer to doctors and patient coordinators. (19 patients, 30 attendants) and 6 experts (4 doctors and 2 patient coordinators). Our findings revealed that patients and attendants appreciated CataractBot because it alleviated hesitations associated with power differences in asking questions, and accommodated individuals with low literacy and tech proficiency through multilingual and multimodal support. Patients and attendants reported that their trust in the system was established through the verification performed by doctors and coordinators. Experts praised the bot as a facilitator, introducing a privacy layer between them and patients, and providing flexibility through asynchronous communication. Moreover, experts noted improved responses over time, likely due to continuous KB updates.
Our main contributions are: (1) The design and development of a novel experts-in-the-loop chatbot222Our code is publicly available at https://github.com/microsoft/BYOeB. , designed in collaboration with doctors and hospital staff, to assist patients undergoing cataract surgery with their information needs. (2) The deployment of our bot in a real-world setting, among 49 cataract surgery patients/attendants at a tertiary eye hospital in India. Four doctors and two patient coordinators served as experts, actively verifying and correcting responses generated by the bot. (3) Drawing from a comprehensive mixed-methods analysis of gathered data, the paper offers pivotal insights specific to CataractBot and broader implications of LLM-powered experts-in-the-loop chatbots.
2. Related Work
Our work is primarily informed by two areas of relevant research: solutions addressing patients’ information needs and the use of conversational agents, including those powered by LLMs, in healthcare settings.
2.1. Patients’ Information Needs
Patients and their attendants seek information throughout the various stages of their medical journey, such as preparing for doctor consultations (Radionova et al., 2023; Leong et al., 2018), verifying diagnoses and treatment plans (Eriksson-Liebon et al., 2021; Bickmore et al., 2009; Christalle et al., 2019), assessing the need for clinical interventions (Yang et al., 2024; Tseng et al., 2023), and aiding in the recovery process (Wang et al., 2021a; Civan-Hartzler et al., 2010; Fadhil, 2018). Studies emphasize that satisfying these diverse information needs improves patient understanding, motivates adherence to treatment regimens, and contributes to overall satisfaction with healthcare services (Tang and Newcomb, 1998). Patients primarily trust doctors with their medical queries (Ding et al., 2020). However, traditional in-person consultations pose difficulties, including high costs (Bodenheimer and Smith, 2013), long wait times, and a strong hierarchy in the patient-physician relationship (Fochsen et al., 2006). Patients often miss opportunities to ask additional (clarification) questions during these brief encounters, as they face information overload (Ferré et al., 2020; Kinnersley et al., 2007). Doctors also struggle with time due to staff shortages straining healthcare ecosystems (Bodenheimer and Smith, 2013; Akerkar and Bichile, 2004). Thus, patients and caregivers are increasingly turning to the Internet for easier access to medical information (Xiao et al., 2023; Chandwani and Kulkarni, 2016). However, the challenge lies in identifying relevant and appropriate information. Individuals struggle to understand complex health information or may even consume inaccurate information, making the process frustrating (Zhang, 2010; Arora et al., 2008; Clarke et al., 2016b).
Information seekers may also consult experienced patients within their social networks who can share valuable insights related to managing similar health conditions (Civan-Hartzler et al., 2010; Skeels et al., 2010). In the absence of such known individuals, people may resort to social media, which is not only ineffective for quality health support due to widespread misinformation and poor management (Malki et al., 2024), but also risky in terms of personal privacy (De Choudhury et al., 2014). Chandwani and Kulkarni (2016) found that although most doctors regard the practice of seeking health information online as an affront to their authority, some see it as inevitable–and even as an opportunity to elevate patients’ health literacy through technology.
Recent work in the field of Computer-Human Interaction has explored novel avenues to address these medical information gaps. For instance, Leong et al. (2018) propose interactive ‘science museum’ exhibits in clinical waiting rooms to educate children and their parents about sickle cell anemia, Wilcox et al. (2010) and Bickmore et al. (2009) propose information displays in hospital rooms to provide patients with real-time status updates and treatment details, and Pfeifer Vardoulakis et al. (2012) propose using mobile phones to present dynamic, interactive reports on patients’ progress and care plans throughout their emergency department stay. Although these are specific solutions for certain diseases and infrastructures, discussions have been ongoing around making Electronic Health Record (EHR) data accessible to patients (Cajander and Grünloh, 2019b; Tang and Lansky, 2005; Chung, 2017). Patient-facing EHR services have received mixed reviews, with healthcare professionals expressing concerns about adding strain to already understaffed healthcare systems, while patients generally welcome the idea (Cajander and Grünloh, 2019b).
In summary, patients need reliable, relevant, and timely information. However, accessing this from busy doctors and hospital staff is challenging. This dichotomy results in patients either receiving generic, unverified, and hard-to-understand information from the internet and social circles or if lucky, obtaining personalized, verified answers, albeit at a higher cost, from healthcare experts. Our solution aims to strike a balance by providing a generic, high-quality answer instantly using LLMs, which is later verified by a medical professional with minimal increase to their workload.
2.2. Chatbots in Healthcare
Chatbots enable natural language conversation, resulting in a minimal learning curve and offering a personalized experience (Jain et al., 2018; Jo et al., 2023; Roller et al., 2020; Ferré et al., 2020; Chung, 2017; Desai and Chin, 2023). They enable users to build complex queries message by message while retaining conversation context (Xiao et al., 2023). Researchers have applied chatbots in various healthcare settings (Wutz et al., 2023), including mental health support (Kumar et al., 2023; Jo et al., 2023; Ma et al., 2023), appointment scheduling (Ittarat et al., 2023), seeking information for chronic disease management (Montagna et al., 2023), and assistance for new mothers (Yadav et al., 2019). Commercial chatbots, such as Babylon Health, Ada, and Florence333Babylon Health: https://www.emed.com/uk. Ada: https://ada.com/. Florence: https://florence.chat/., allow individuals to inquire about their health issues. All of these bots are fully automated, with some even using LLMs. However, the involvement of doctors in these bots is limited to the data curation step (Xiao et al., 2023). Consequently, they are susceptible to generating inaccurate information, rendering them “not ready for clinical use” (Au Yeung et al., 2023).
LLMs have been criticized for hallucinating and not displaying the reason behind their outputs (Fogel and Kvedar, 2018). Several methods have been proposed to address these issues, including chain-of-thought prompting (Wei et al., 2024), retrieval-augmented generation (RAG) (Lewis et al., 2020), and few-shot learning. Prior work suggests that healthcare professionals are more likely to adopt AI tools when they perceive them as ‘partners’ that enhance their workflow and expertise (Henry et al., 2022). Seitz et al. (2022) identified that patients’ trust in healthcare chatbots is primarily based on a rational evaluation of their capabilities, while trust in human medical professionals relies more on emotional connections and personal rapport. Integrating doctors into the loop with automated chatbots could potentially combine the best of both worlds, a concept we propose in CataractBot.
Given our specific focus on cataract surgery, we also examined literature at the intersection of ophthalmology and LLMs. Bernstein et al. (2023) reported that responses generated by LLMs to patient eye-care-related questions are comparable to those written by ophthalmologists. Based on similar findings, Ittarat et al. (2023) proposed integrating chatbots into ophthalmology practices to provide 24/7 support for common inquiries on eye conditions. Additionally, Yilmaz and Dogan (2024) reported that chatbots are an effective tool for educating cataract surgery patients. Our work contributes to this body of literature.
Beyond chatbots, prior work has explored “chat” as an interface to connect patients and doctors in both synchronous (Li et al., 2023) and asynchronous (Johansson et al., 2020a, b) manners. Synchronous messaging requires both the patient and the doctor to be simultaneously active, allowing instant feedback and coherent dialogue. On the other hand, asynchronous messaging platforms like WhatsApp (Yadav et al., 2019), WeChat (Wang et al., 2020), SMS (Perrier et al., 2015), email (Makarem and Antoun, 2016), or web portals (Mamykina et al., 2008; Tuli et al., 2018) enable communication at the convenience of both parties, providing flexibility. Previous research indicates that both doctors and patients prefer asynchronous texts over instant audio/video calls (Stamenova et al., 2020; Johansson et al., 2020a), despite lacking instant feedback (Li et al., 2023). In our work, we combine both approaches—leveraging LLM-based responses in the synchronous mode and having a human medical expert respond in the asynchronous mode. This hybrid approach represents a novel and effective method for addressing information needs in the healthcare domain.
Finally, the use of AI in real-world healthcare deployments is limited due to socio-technical uncertainties in the last mile (Osman Andersen et al., 2021). To avoid such shortcomings, as recommended by Thieme et al. (2023b, a), we designed and studied our system within a tertiary eye hospital in India and validated its utility to understand its integration into clinical workflows.
3. Formative Study
The development of CataractBot was informed by our literature survey, our understanding of the Indian healthcare ecosystem, and insights from a formative study conducted at Sankara Eye Hospital. To identify key chatbot requirements, we conducted semi-structured interviews with 4 ophthalmologists and 2 patient coordinators between July-Aug 2023. From these interviews, we aimed to understand their current workflow, the types of questions patients ask, and their concerns regarding patients’ knowledge gaps. Participants had an average age of 38.58.8 years, with 11.07.5 years of work experience. They participated in the study voluntarily without compensation. The interviews were conducted in English and lasted around 45 minutes each. All sessions were audio-recorded and transcribed by the interviewer. We conducted a thematic analysis to identify these key design requirements for the chatbot.
Accuracy: The chatbot’s ability to answer both medical and logistical questions with accuracy, precision and contextual relevance was found to be essential given the medical context. While a standard LLM like GPT-4 (OpenAI et al., 2023) might offer accurate information, it may not be tailored to India- or hospital-specific nuances. For instance, doctors informed us that although the phacoemulsification method is prevalent globally for cataract surgeries, they often favor manual small incision-based procedures (Singh et al., 2017) due to their cost-effectiveness. Similarly, cultural considerations significantly influence pre- and post-surgery questions. As a doctor stated, “People will have weird questions… like ‘when can I do ayurvedic nasal therapy post-op?’.” Questions related to diet require awareness of Indian food options for accurate guidance. To ensure accurate and contextual responses, our chatbot relies on a Knowledge Base (KB) curated by hospital medical staff, including doctors, patient coordinators, and members of Quality Control and Patient Safety team.
Trustworthiness: Access to trustworthy information is crucial during medical treatment. As one doctor noted, “The success of treatment is not only in medicine, but the trust they place on us.” Patients and their attendants typically place their utmost trust in the operating doctor and hospital staff. Even when patients learn something from the internet or friends, they often confirm that information with their doctor. E.g., a doctor noted, “They come and ask us… my friends have told me this… is it correct or not?” To ensure trustworthiness in CataractBot we decided that (a) each response will be verified and, if necessary, edited by doctors and patient coordinators, and (b) patients (or their attendants) will be explicitly notified when this expert verification happens. Note: We intentionally designed CataractBot to be non-specific to individual patients by not connecting it with the hospital’s scheduling system or EHR data. This decision was made because LLMs can memorize training data, including personally identifiable information (PII) like emails and phone numbers, and leak it during inference (Borkar, 2023), raising privacy concerns and potentially breaking users’ trust.
Timeliness: Discussions with hospital staff highlighted the crucial need for real-time responses. Patient coordinators noted that patients frequently reach out to them over multiple phone calls to ask logistical questions and relay medical queries to their consulting doctors. Due to busy schedules of doctors and coordinators, immediate responses are not always feasible. This often results in patient seeking information from other sources (including friends and online sources), which can propagate misinformation, or fail to completely address their information needs. To counter this, we decided that the chatbot should provide an instantaneous response to every patient query using a custom KB curated by hospital staff. Subsequently, these responses gets reviewed and refined by doctors/coordinators. Moreover, doctors mentioned their unavailability for several hours while performing surgeries. To ensure swift verification of every query, CataractBot forwards queries to an ‘escalation’ expert if they remains unverified for 3 hours.
Usability: After cataract surgery, patients are expected to reference their discharge summaries for any queries, but this can be challenging due to their temporarily reduced vision. Prior research (Grönvall and Verdezoto, 2013) has identified the crucial role immediate family members play in monitoring the patient’s healthcare. Coupled with doctors’ input (“Hardly 10% of questions are asked by patients; attendants are the one who ask.”), this underscored the necessity for the bot to be usable by the four key stakeholders of the cataract surgery ecosystem: patients, their attendants, doctors, and patient coordinators. To cater to diverse backgrounds, literacy levels, and varying degrees of tech-literacy among hospital visitors, as suggested by doctors and patient coordinators, we made the chatbot multilingual and multimodal similar to (Jain et al., 2018; Findlater et al., 2009). We chose WhatsApp as the messaging platform due to its widespread use in India (Statista, 2022), despite the richer feature set of Telegram or the development flexibility of a custom chatbot. This choice aims to minimize the learning curve and ensure quick onboarding without requiring additional app installations. Considering doctors’ demanding schedules, often involving consulting 50 patients and performing 10 cataract surgeries a day, we implemented features to minimize their bot-related workload: (1) one-click interaction to verify an answer, (2) allow experts to provide corrections using informal and ‘symbolic’ messages (Beasley, 2009), (3) handle spelling and grammar errors, and (4) use expert edits to enhance the bot’s KB, minimizing similar edits in future.
4. CataractBot System Design
Here, we describe the CataractBot system (Figure 3), including its key components–input language and modality, response generation and verification, and escalation and reminders–and provide a detailed account of its implementation.
4.1. Components
4.1.1. Input Language and Modality
Sankara Eye Hospital in Bangalore (in the state of Karnataka), referred to as the Silicon Valley of India, caters to patients from various linguistic, educational, and technical backgrounds, including those from the Information Technology sector and neighboring states. Analysis of a 2011 census highlights Bangalore as one of India’s most linguistically diverse cities (Office of the Registrar General & Census Commissioner, 2011). To accommodate this diverse patient demography, CataractBot is designed to support five languages: English, Hindi (the local language of nine states in India), Kannada (the local language of Karnataka), and Tamil and Telugu (the local languages of two neighboring states of Karnataka). As the expert may not be proficient in all five languages, their interactions with the bot are exclusively in English. Upon onboarding, the language preference of the patient and their attendant is collected through an online form, and CataractBot initiates the conversation by sending a set of ‘welcome messages’ in the chosen language. Among these messages is an option allowing users to modify their preferred language in future. To accommodate different levels of literacy, including literate, semi-literate, and illiterate users, CataractBot supports both text and speech inputs. For every voice message, CataractBot sends both a text response and a spoken version of the text as an audio message.
Recognizing the challenge users face in initiating a conversation, often referred to as ‘Blank Page Syndrome’ (Moden, 2022; YM et al., 2023), CataractBot provides a set of three frequently asked questions along with the welcome message. Additionally, every bot’s response includes a “What to do next?” prompt, offering three related questions based on the preceding query (Figure 2F). This feature lowers the barrier to entry (as formulating questions has been found to be non-trivial for low-literate individuals (Tseng et al., 2023)), enhances the natural flow of conversation, and saves the user time.
4.1.2. Response Generation, Verification and Icons
Upon receiving a message, CataractBot classifies it as a medical question (e.g., dos and donts before and after surgery), logistical question (e.g., scheduling or insurance related), small talk, or ‘other’, and provides a response in real-time.
For medical and logistical questions, the bot strictly employs the KB curated by the hospital’s medical team to generate an appropriate response.
The response includes a question mark icon
as a ‘reaction’ indicating the unverified status (Figure 2E).
In instances where the custom KB lacks an answer, the bot responds with a template “I don’t know” response.
For small talk messages (such as “Hello” or “Thank you for the information”), the chatbot provides corresponding small talk responses.
For medical questions, the operating doctor of the respective patient receives a message comprising of the question asked, the bot’s response, and patient demographics (Figure 1A).
As Rajashekar et al. (2024) found that citations improve trust in LLM-generated responses, this message also includes citations of the documents used to generate the response.
The doctor is prompted with the question, “Is the answer accurate and complete?” offering three response options: Yes, No, and Send to Patient Coordinator (Figure 1A).
Tapping ‘Yes’ replaces the question mark icon
in the information seeker’s received answer with a green tick icon
, confirming verification (Figure 1F). Additionally, the bot notifies the information seeker that the answer has been verified, tagging that particular response (Figure 1G).
Tapping ‘No’ replaces the question mark
with a red cross icon
, indicating an incorrect answer (Figure 1C), and asks the information seeker to await a corrected response from the doctor (Figure 1D).
The doctor is asked to provide a correction; they are not required to edit the bot’s message, instead, they can offer a correction in free form text.
The CataractBot system automatically combines the bot’s initial answer with the doctor-provided correction to frame a new response, which is delivered to the user with a green tick icon
(Figure 1E).
This new response is not sent to the doctor, to minimize their workload by limiting verification to a single round.
If a question is misclassified, i.e., a logistical question is sent to the doctor, they have the option to ‘Send to Patient Coordinator’.
Similar to the doctor’s workflow, the patient coordinators verify and provide corrections to the bot’s response for logistical questions.
We deliberated between displaying unverified responses in real-time and presenting only verified responses post-expert verification and correction.
Through our
formative study and discussions with hospital staff, it became clear that providing real-time responses was crucial.
Delayed responses might lead to patients making repeated calls to the hospital or resorting to online sources, failing to effectively address their information needs.
4.1.3. Escalation Mechanism
The CataractBot system employs an intricate escalation and reminder mechanism to ensure swift verification.
In the doctor’s WhatsApp interface, unanswered questions are marked with a question mark icon
(Figure 2E) and answered questions with a green tick
, enabling them to easily identify pending queries.
If an answer is not verified by the operating doctor within three hours, it is automatically sent to the designated escalation doctor.
The operating doctor receives a notification about the escalation.
Both the operating and escalation doctors can then verify/correct the response.
Selecting ‘Yes’ or ‘Send to Patient Coordinator’ immediately marks the query with a green tick (Figure 1F) for both the doctors, indicating task completion.
However, if either doctor selects ‘No’, the green tick appears only after that doctor provides a correction.
If neither the operating nor escalation doctor verifies a question within six hours, both receive a reminder notification indicating the pending status.
Additionally, every four hours during working hours (at 8 am, 12 pm, and 4 pm), a list of all questions pending verification for more than six hours is sent to both the operating and escalation doctors.
This additional reminder was added based on the doctors’ feedback.
This workflow is mirrored for the patient coordinator and escalation patient coordinator.
4.1.4. KB Update Process
To minimize experts’ labor, we use the expert-provided edits to update the KB, which increases the likelihood of ‘Yes’ responses from experts for similar questions in the future. However, certain responses, such as those specific to individual patients (e.g., “When to reach the hospital for my surgery?”) must not be added to the KB as they are not generalizable. We appointed a senior cataract surgeon as the ‘KB update expert’, who received a spreadsheet via email at 8pm daily. Their task involves reviewing each row and determining whether the information should be added to the KB (by responding with a ‘Yes’/‘No’ in the ‘Should Update KB?’ column), and if so, modifying as needed the ‘Final Answer for KB’ column, containing the bot’s updated answer based on the expert’s correction. At 3 am daily, the system extracts ‘Question’ and ‘Final Answer for KB’ data from all rows marked ‘Yes’, and append these to the KB in the ‘expert-FAQ’ document. Additionally, we prompt the LLM to prioritize this ‘expert-FAQ’ document within the KB. This ensures that CataractBot improves with these updates, gradually enhancing its accuracy over time.
4.2. Implementation Details
The CataractBot system relies on these five components (Figure 3): (a) LLM: for response generation, (b) Vector Database: for storing and retrieving the custom KB, (c) Language Technologies: for translation and transcription, (d) WhatsApp Services: for message exchange, and (e) Cloud Storage: for storing interaction logs.
We opted for GPT-4, the leading LLM at the time of our system development in May 2023 (OpenAI et al., 2023). The custom KB comprises of various cataract surgery and hospital-specific documents, including Consent Form, SOP (Standard Operating Procedure), FAQs, Pre- and Post-Operative Guidelines, and Hospital Information. These documents are ingested as data chunks into a Chroma vector database, using the OpenAI model ‘text-embedding-ada-002’ to generate embeddings. Upon receiving a question, GPT-4 classifies the query type. For a medical/logistical question, CataractBot employs a RAG (Lewis et al., 2020) approach. This involves performing a vector search on the KB to extract the three most relevant data chunks related to the query. GPT-4 is then prompted (full prompt available in Appendix A) to generate an answer for the query from these data chunks. If the answer is not present, the bot responds with an “I don’t know” message.
The prompt underwent several iterations. To validate its effectiveness, a doctor and a patient coordinatorrecorded 153 questions posed by patients/attendants scheduled for cataract surgery over a week. We used our GPT-4 prompt to obtain answers to these questions from the custom KB, and subsequently the generated responses were evaluated by the same doctor and coordinator. This not only helped refine the LLM prompt, but also led the medical staff to improve existing documents and contribute additional documents to enrich the KB. This process was iterated thrice.
Additionally, we utilize GPT-4 to incorporate corrections provided by doctors/coordinators to generate the final expert corrected response. This involved prompting GPT-4 with the patient’s question, the initial response from the bot, and the expert-provided correction, and asking it to generate a revised answer considering the expert’s input. Our experience showed that GPT-4 executed this task well. Regarding language integration, GPT-4 primarily comprehends and responds in English (Ahuja et al., 2023). To bridge the language gap, our system adhered to the standard approach (ait, [n. d.]) of translating Indic languages into English for input using Azure AI Translator. It then translated GPT-4’s English responses back into the respective Indic languages for output. To facilitate speech as input and audio as output, CataractBot uses Azure speech-to-text and text-to-speech models. (Note: GPT-4 lacks support for speech input.) Finally, all text and audio interactions between the CataractBot system and users (including patients, attendants, operating and escalation doctors, coordinators, and KB update expert) get logged in a cloud storage for further analysis.
5. Study Design
To evaluate the effectiveness of this LLM-powered experts-in-the-loop chatbot in addressing the informational needs of patients undergoing cataract surgery, we conducted a mixed-method user study during Nov 2023-Jan 2024 at our collaborators’ institute, Sankara Eye Hospital. This is one of the leading tertiary eye care and teaching institutions in Bangalore, India. It typically attends to more than 500 patients every day, and conducts over 50 cataract surgeries daily. This study was approved by both the Scientific and Ethics Committees of Sankara Eye Hospital. None of the study participants—patients, attendants, doctors, or coordinators—received any financial incentives for their participation, in accordance with hospital norms.
5.1. Procedure for Patients and Attendants
As per the hospital’s protocol, once a patient opts for cataract surgery based on their doctor’s recommendation, the patient and their attendant needs to meet a patient coordinator. The coordinator helps with general guidance on pre- and post-operative measures, scheduling of the surgery date, and logistical arrangements. Toward the end of this interaction, the coordinator was tasked to assess the patient’s eligibility for the study based on specific criteria: (a) age of 18 or older, (b) fluent in one of the five languages supported by CataractBot, (c) scheduled for cataract surgery within a week, (d) no history of cataract surgery in the last 6 months (as recent patients would likely have minimal informational needs), and (e) having one of the three participating operating doctors as their surgeon. If these criteria were met, the patient and their attendant were directed by the coordinator to meet a researcher (the first author) stationed at the hospital.
The researcher introduced them to CataractBot and outlined the study’s protocol which involved using the bot for around two weeks, engaging in an interview pre and post-surgery, and completing a chatbot usability form. If they agreed to participate, participants were requested to sign a consent form, after which the researcher filled a web-based onboarding form. This form included details such as the patient and attendant’s phone numbers linked to WhatsApp, preferred languages, consulting doctor and coordinator, surgery date, and basic demographic information (age, gender, and education). Upon form submission, participants received ‘welcome messages’ from CataractBot. The researcher instructed them to ask a trial question, either by choosing from suggested questions or by typing/speaking in their preferred language. Upon receiving a response, participants were briefed on the icons and expert verification system. Throughout this process, the researcher encouraged them to ask any questions regarding the bot’s usage. After onboarding, participants received two daily reminder messages (at 7:30 am and 4 pm) prompting them to “Feel free to ask any cataract surgery related questions” until one week post surgery, when their access was automatically revoked.
We conducted two semi-structured interviews with both the patient and their attendant: one on the surgery day and another a week post-surgery. These specific days were chosen because they coincided with the patient’s required hospital visits. The interviews explored their overall experience, specific features they liked or disliked, suggestions for improvement, and questions regarding the trustworthiness, timeliness, and accuracy of responses. Both interviews followed a similar structure, differing only in how the bot supported the participant before versus after surgery. After each interview, participants were requested to fill out a chatbot usability form. They rated their chatbot experience through eight questions, using a five-point Likert scale, covering aspects like ease of use, understandability, and timeliness (Appendix A.2). These questions were adapted from metrics outlined in prior usability evaluations of healthcare bots; previous work used scales that included aspects not relevant to our study, such as the chatbot’s visual design (Abd-Alrazaq et al., 2020; Holmes et al., 2023). Patient-attendant pairs were interviewed together but were asked to fill separate usability forms.
All interviews with patients and attendants were conducted by the first author in English, Hindi, Kannada, or Tamil, as the researcher is fluent in these languages. (Note: Participants who selected Telugu as their preferred language during onboarding were interviewed in one of the other four languages.) Interviews on the day of surgery took place in person at the hospital, while seven days post-surgery interviews were conducted either in person at the hospital or through pre-scheduled telephone calls. This decision was influenced by the substantial waiting time on the surgery day—patients typically spend about five hours in the hospital, with the procedure itself taking 45 minutes. Conversely, during the post-surgery visit, patients receive priority and experience minimal waiting time. Each interview typically spanned 20 to 40 minutes and was audio-recorded with the participant’s consent.
5.2. Procedure for Doctors and Coordinators
In addition to patients and attendants, we recruited four doctors and two patient coordinators. Among the doctors, three served as operating doctors, and one as both an escalation doctor and KB update expert. Both patient coordinators specialize in handling cataract surgery patients, with one serving as a coordinator and the other as an escalation coordinator. The escalation doctor and escalation coordinator held senior positions in the hospital and carried administrative responsibilities. To ensure preparedness, all participating experts attended one of two one-hour training sessions conducted by three researchers. During these sessions, CataractBot was demonstrated, and they were onboarded onto the platform. The researchers assumed patient roles, asking medical and logistical questions for the doctors and coordinators to verify and edit. The first session revealed a few bugs, and feedback was collected to enhance the bot. For example, a doctor suggested displaying the question first (rather than the patient’s demographics) in verification messages, as only the top portion is visible in ‘tagged’ WhatsApp messages (Figure 1E). After using CataractBot for two weeks or longer, the participating experts were interviewed by the first author. These interviews followed a semi-structured format, focusing on the bot’s usability, its integration into their workflow, and its impact on interactions with patients. The interviews were audio-recorded, conducted in English, in-person, and typically lasted 45-60 minutes.
5.3. Participants
A total of 31 patient-attendant pairs (Table 1), consisting of 19 patients and 30 attendants, took part in our study. Due to the predominantly elderly demographic undergoing cataract surgery, 12 patients lacked access to WhatsApp/smartphones and were unable to participate but their attendants used the bot. One patient was not accompanied by an attendant. The average age of the 19 participating patients (3 females) was 58.88.01 years. Although females comprised 32.7% of our study participants, only 15.8% of the patients were female. This aligns with prior work showing a gender disparity in accessing cataract surgery in India (Prasad et al., 2020; Ye et al., 2020). Patients had diverse preferred languages (5 English, 5 Hindi, 5 Kannada, 3 Tamil, and 1 Telugu) and education levels (6 Grade 10, 4 Grade 12, 6 Bachelors, and 3 Masters). In contrast, the attendants were younger, predominantly fluent in English, and well-educated. The average age of the 30 participating attendants (11 female) was 37.210.3 years. A majority (23) preferred English, with 2 each favoring Hindi, Kannada, and Tamil, and 1 preferring Telugu. Education levels were mostly high, with 10 having Masters and 13 having Bachelors degrees. Among the 31 patient-attendant pairs, 22 participated in the surgery-day interview, and 10 took part in the one-week post-surgery interview. Various factors influenced participation, including time constraints (resulting in a few declined interviews) and data reaching saturation. Additionally, four doctors and two patient coordinators participated in the study (Table 2). Note: There was a partial overlap between participants in this study and the formative study, with two doctors and one coordinator participating in both.
| PId | Age | Sex | Language |
AId | Age | Sex | Language |
Days |
| A1 | 43 | M | English | Masters | Yes | Yes | 7 | 9 | |||||||||||||||
| A2 | 31 | F | Kannada | Grade 12 | Yes | Yes | 3 | 13 | |||||||||||||||
| P3 | 56 | M | Tamil | Grade 10 | A3 | 35 | M | English | Bachelors | Yes | No | 4 | 6 | ||||||||||
| P4 | 52 | M | English | Grade 12 | A4 | 39 | F | English | Grade 12 | Yes | No | 4 | 8 | ||||||||||
| P5 | 67 | M | Kannada | Grade 10 | A5 | 40 | M | Kannada | Bachelors | Yes | No | 5 | 4 | ||||||||||
| A6 | 35 | M | Telugu | Grade 10 | Yes | No | 1 | 2 | |||||||||||||||
| P7 | 58 | M | Telugu | Grade 10 | A7 | 25 | M | English | Masters | Yes | Yes | 23 | 17 | ||||||||||
| A8 | 39 | M | Hindi | Bachelors | Yes | No | 6 | 6 | |||||||||||||||
| P9 | 69 | M | English | Masters | A9 | 38 | F | English | Masters | Yes | Yes | 6 | 14 | ||||||||||
| A10 | 71 | M | English | Bachelors | Yes | Yes | 26 | 15 | |||||||||||||||
| A11 | 32 | M | Hindi | Grade 10 | Yes | No | 14 | 6 | |||||||||||||||
| P12 | 65 | M | Hindi | Bachelors | A12 | 31 | F | English | Masters | Yes | Yes | 2 | 6 | ||||||||||
| P13 | 69 | M | Tamil | Grade 12 | A13 | 35 | M | English | Masters | Yes | Yes | 25 | 18 | ||||||||||
| P14 | 56 | F | Kannada | Grade 10 | A14 | 33 | F | English | Bachelors | Yes | Yes | 22 | 30 | ||||||||||
| A15 | 57 | M | English | Grade 12 | Yes | No | 4 | 4 | |||||||||||||||
| P16 | 57 | F | Hindi | Grade 10 | A16 | 25 | M | English | Masters | Yes | No | 1 | 1 | ||||||||||
| A17 | 39 | M | English | Bachelors | Yes | Yes | 2 | 8 | |||||||||||||||
| P18 | 56 | M | English | Grade 12 | A18 | 23 | M | English | Masters | Yes | Yes | 27 | 24 | ||||||||||
| P19 | 41 | M | English | Masters | A19 | 36 | F | English | Bachelors | Yes | No | 19 | 15 | ||||||||||
| P20 | 64 | M | Tamil | Grade 12 | A20 | 60 | F | Tamil | Grade 12 | Yes | No | 4 | 16 | ||||||||||
| A21 | 43 | M | English | Masters | Yes | No | 15 | 19 | |||||||||||||||
| A22 | 35 | F | English | Bachelors | Yes | No | 7 | 11 | |||||||||||||||
| P23 | 54 | M | English | Bachelors | 7 | 8 | |||||||||||||||||
| P24 | 66 | M | Kannada | Bachelors | A24 | 37 | M | English | Masters | No | No | 1 | 1 | ||||||||||
| A25 | 40 | F | Tamil | Grade 10 | No | No | 7 | 3 | |||||||||||||||
| P26 | 65 | M | Hindi | Bachelors | A26 | 34 | M | English | Bachelors | No | No | 19 | 15 | ||||||||||
| P27 | 64 | M | Hindi | Bachelors | A27 | 35 | M | English | Bachelors | No | No | 18 | 14 | ||||||||||
| A28 | 40 | M | English | Masters | No | No | 1 | 1 | |||||||||||||||
| P29 | 45 | F | Kannada | Grade 10 | A29 | 29 | F | English | Bachelors | No | No | 15 | 10 | ||||||||||
| P30 | 50 | M | Hindi | Masters | A30 | 25 | F | English | Bachelors | No | No | 5 | 3 | ||||||||||
| P31 | 63 | M | Kannada | Bachelors | A31 | 32 | M | English | Bachelors | No | No | 15 | 36 |
| Id | Role | Age | Sex | Languages | Highest Education | Experience |
| D1 | Operating Doctor | 42 | F | English, Hindi, Kannada, Tamil, Telugu | Masters | 10+ years |
| D2 | Operating Doctor | 44 | F | English, Hindi, Kannada, Tamil, Telugu, Malyalam | Masters | 15+ years |
| D3 | Operating Doctor | 46 | M | English, Hindi, Kannada, Tamil | Masters | 20+ years |
| D4 | Escalation Doctor; KB update expert | 46 | M | English, Hindi, Kannada, Tamil | Masters | 20+ years |
| C1 | Coordinator | 33 | F | English, Hindi, Kannada, Tamil, Telugu | Bachelors | 10+ years |
| C2 | Escalation Coordinator | 36 | F | English, Hindi, Kannada, Tamil, Malyalam | Bachelors | 15+ years |
5.4. Data collection and analysis
We performed a mixed-method analysis to systematically analyze the collected data comprising of usage logs, interview transcripts, and usability survey responses. CataractBot interaction logs and usability survey data were quantitatively analyzed. Descriptive statistics and statistical tests (such as t-tests and ANOVAs) were used to analyze the count and type of questions asked, verification styles and response times of doctors and coordinators, edits performed by the KB update expert, and responses to the usability survey. Note: For ANOVAs, the sphericity assumption was tested using Mauchly’s test, and in case of sphericity violations, the Greenhouse-Geisser correction was applied.
The first author translated and transcribed all interviews (totalling 11.5 hours) into English soon after they were conducted. Throughout the study, all authors regularly engaged in discussions to review observations and interviews. To analyze interview transcripts, we employed an inductive coding approach, with the first author open-coding (Williams and Moser, 2019) all interviews line-by-line. Subsequently, two authors collaboratively reviewed these codes to identify interesting themes in the data (similar to (Okolo et al., 2021)). We iterated on these themes to distill higher-level themes (like “Usage”, “Features”, and “Privacy and Trust”) that we present in our findings below. Note: While there were a total of 49 patients and attendants using CataractBot, at times, we treat them as 31 participants (patient-attendant pairs). This choice stems from instances highlighted during interviews, where patients instructed attendants to query the bot using their phones. Additionally, both were interviewed together. Only when comparing their distinct roles, we analyzed their usage log data separately.
5.5. Positionality
All authors are of Indian origin, and currently reside in Bangalore, India. Among them, two identify as female and the rest as male. One author is a practising ophthalmologist, performing over 10 cataract surgeries weekly, and another author is part of the Quality Control and Patient Safety team at the same hospital. Both were involved in designing the study, shaping research questions, providing feedback on initial bot versions, curating the KB, and ensuring that any risk of harm to the patients was mitigated. The ophthalmologist also served as the KB update expert. Two authors specialize in HCI and design, three possess healthcare research experience, and four have expertise in software development. We approached this research drawing on our individual experiences and learning from working at the intersection of computing and healthcare in the Indian context. Our interest in making healthcare more accessible to the masses has informed the design of CataractBot, guided our study design, and shaped our data analysis.
6. Findings
Overall, 31 patient-attendant pairs sent 343 messages to CataractBot, with an average of 11.068.44 messages per pair. Out of the 343 messages, 225 were medical questions, 87 logistical, 27 small talk, and 4 others. Doctors directly approved (said ‘Yes’ to) 69.8% (157) of the bot-generated answers to medical questions, while patient coordinators directly approved 58.6% (51) of logistical answers.In the usability survey, 77.3% of participants found CataractBot’s responses to be useful, appropriate, and informative, while 91.0% said they would be willing to use it in the future. Below, in our findings, we delve into how participants used CataractBot (answering RQ1 and RQ3), examine the impact of novel features introduced in CataractBot (answering RQ2), and explore privacy and trust implications of using an LLM-powered chatbot (answering RQ1).
6.1. CataractBot Usage
6.1.1. Reasons for (Not) Using CataractBot
Information seekers used CataractBot to address questions they had forgotten or were uncomfortable to ask the doctor or patient coordinator, clarifying answers, and seeking updates. Doctors stated that most information seekers now-a-days come prepared for face-to-face conversations, optimizing their limited time with experts and reducing multiple visits. CataractBot can alleviate the burden of remembering every question. Ten participants highlighted that the bot helped in recalling information provided by the doctor/coordinator during in-person consultations, which was either forgotten or incompletely understood. As A13 stated: “We are humans… cannot give complete attention all the time.” Patients found comfort in “written” information, ensuring persistence and enabling them to refer back to it when needed (similar to (Jain et al., 2018)), “I need not remember everything.” (A20).
Due to existing power differences between doctors and patients (Joseph-Williams et al., 2014), individuals often hesitate to ask questions. Experts acknowledged that educated individuals have access to other information sources, and are comfortable posing questions. In contrast, “less educated, local language speaking patients” are the ones for whom CataractBot “can do wonders” (D3). Five participants, including three with Grade-12 education, exclusively relied on CataractBot for information (asking 57 questions in total) as no one in their social circles had experience with cataract surgery, and they were not accustomed to seeking information online. Another expressed that checking with the bot was more comfortable than approaching a doctor, citing the bot’s non-judgemental nature: “(The patient) has cholesterol issues and heart surgery… the doctor would make faces when I ask about food intake… but not the bot.” (A4). Six participants also mentioned using the bot to “double-check” (A7) and seek “reassurance” (A17), in spite of knowing certain answers.
The asynchronous feature was helpful, as the information seekers did not have to constantly wait, either physically or over a call, for an expert’s response. Seven participants emphasized that the bot saved them time, as it eliminated the need to visit or call the hospital. Traveling to the hospital can be time-consuming, and contacting busy hospitals over the phone is challenging due to high patient loads, as “the phone is always busy” (A4). Also, information seekers found the bot easy to use, with 86.2% rating ‘Agree’ or ‘Strongly Agree’ in response to the usability question (Appendix A.2), mainly due to their familiarity with WhatsApp. Additionally, features such as support for multiple languages, audio support, and related questions (Figure 2), lowered barriers to initiating conversation.
Experts used CataractBot driven by their commitment to help patients.
“I want to help my patients… I am least bothered about the timing. If he is having a problem, he should be answered… after all, we are living for the patient’s sake.” – D2.
Experts also appreciated the intermediary role of the bot, which introduced a layer of separation from patients: “From a doctor’s standpoint, it is very good because it puts in a curtain between us.” (D4).
Despite the compelling reasons to use the bot, out of the 49 information seekers, 8 did not use the bot sending two or less messages. Six participants mentioned that their in-person interactions with the doctor and patient coordinator sufficed to address their queries, reducing their reliance on the bot. Two participants considered cataract surgery to be a “simple” procedure, leading them to have no questions. This perception stemmed from their previous experience with cataract surgery of their other eye. This suggests that CataractBot could play a more pivotal role in primary healthcare centers, where the human infrastructure may be less capable of addressing patients’ information needs or where power imbalances make individuals less likely to seek answers.
Participants also complained about receiving “I don’t know” (IDK) responses (57 total) and experiencing delays in experts verifying responses. A total of 36 (10.5%) messages requested status updates—seeking information about the surgery time and discharge time. Due to privacy concerns, CataractBot was not integrated with the hospital’s patient management system, resulting in IDK responses for such questions. The patient coordinator later provided appropriate responses. While participants acknowledged the appropriateness of IDK responses due to the lack of access to patient details, they suggested providing a “tentative response” if possible. E.g., for the question “Mine is general anaesthesia or local anaesthesia?” (A10), the bot provided an IDK response. However, A10 expected a response such as “99% of cataract surgeries at Sankara Eye Hospital are done under local anesthesia”. From an expert’s perspective, responding to certain IDK questions was challenging as they lacked the full context. E.g., A21 asked: “I want to postpone the surgery to February… is this okay or do I need to get this operated immediately?” In response, the doctor corrected the bot’s IDK response with: “You come and check”. This correction frustrated the attendant, who stated: “But (the patient) already went through the complete eye care check; the doctor know her condition… I don’t want to come and visit the doctor [again]… it will waste 2-3 hours.” (A21). For time-sensitive questions, A1 and A9 desired “quick response” from the bot. The delay in expert verification led them to perceive the bot as ineffective.
“The dosage… every 2 hours take 1 drop. If it takes anything more than 2 hours to respond [to my eye drop related question], the purpose is already defeated… If it is not providing me timely help… as a consumer, I don’t see value add.” – A1.
6.1.2. Workflow
Out of the total 343 messages sent by patients and attendants, all, except four, were sent between 7 am and 11 pm, with an average of 14.313.7 messages/hour. The messages were distributed throughout the day, as A18 mentioned: “Whenever I am idle at work or at home… whatever (question) comes to my mind, I just ask.” The workflow employed by participants varied, with time-sensitivity being a crucial factor. A majority of participants (86.4%) agreed that the bot responded quickly. On average, the bot took 11.810.2 seconds to respond. Expert verification took 162.9172.3 minutes. We found verification without edits (151.1147.1 minutes) to be significantly faster than verification with edits (191.3172.4 minutes) with t(265)=2.7, p¡0.05. For time-insensitive questions, participants were “okay to wait” for the expert’s verification, since the response was being verified by “busy doctors”. This demonstrated their clear understanding of the bot’s workflow.
“I know that it will take its own time… as it’s doctors… based on their busy schedule, their availability… It’s not like there’s some dedicated doctor monitoring it all the time, right? If that is the case, then bot is not required. They can directly reply.” – A17.
A few participants mentioned that they hardly noticed the wait as they were occupied with their day-to-day activities. They would ask a question, leave their phones, and check the answer later when free. In such cases, notifications played a key role as the notification message (Figure 1G), stating that the expert had provided a verified response, made the otherwise unnoticeable green tick (Figure 1F) prominent and the bot’s process self-explanatory.
If participants received an IDK response, they sought information from other sources, including Google and asking friends and family. For time-sensitive questions, participants reported using Google to “double-check” the initial LLM-generated, unverified response due to the delay in obtaining an expert verified response. Four participants, accustomed to relying on Google for their informational needs, reserved CataractBot exclusively for questions requiring expert opinions, such as “When can i do exercises as anulom vilom?” (P27). According to D4, without CataractBot, patients might have resorted to “Google” and risked following inaccurate advice, given the unavailability of doctors for immediate consultations. CataractBot improves the likelihood of receiving accurate responses, as it utilizes a custom KB.
We found that the experts were “fine” with the additional task of verifying the bot’s responses, as articulated by D4: “It takes only about 15-20 seconds to answer a question… It was not really eating into my time… It’s much easier than replying to my mailbox.” The majority of responses (31, 10.8%) were provided by experts at 4-5 pm, post their busy workday, as most OPD and surgeries finish by 4 pm. The second-highest responses (28, 9.8%) occurred at 1-2 pm, during their lunch break. This usage pattern was described as:
“I use my mobile phone only when I go for breaks… lunch break, coffee break, or in the morning before I start OPD. The rest of the time, my (cellular) data is off.” – D1.
As five experts used CataractBot on their personal WhatsApp accounts, it minimized the learning curve but also blurred the boundary between their home and work, similar to previous findings (Mols and Pridmore, 2021; Wang et al., 2020). Our experts also noted that WhatsApp had a low signal-to-noise ratio, resulting in instances where they missed CataractBot messages: “I am not much into WhatsApp. I just see messages in the important groups. Otherwise my WhatsApp has 5000+ [unread] messages… college groups, family groups.” (D2). Addressing this issue, the escalation doctor D4 suggested that experts should be recommended to “pin” CataractBot on WhatsApp, ensuring it remains consistently at the top of their chat list.
6.1.3. Question Type and Style
To analyze the medical and logistical questions posed by participants, one of the researchers manually classified them into 12 broad categories, identified in a bottom-up manner (similar to (Li et al., 2023; Clarke et al., 2016b)). Note: Several questions fell into multiple categories. Figure 4A illustrates the relative significance of these categories before, on, and after the day of surgery. Prior to the surgery, patients and attendants were concerned about the surgical procedure (asking questions about “anaesthesia”, “lens”, and “duration of surgery”) and sought guidance on how to prepare for it (e.g., “eat or not before the surgery”, “documents to bring for surgery”, and “take blood thinners on the day of surgery”). One day before the surgery, questions related to the “schedule” peaked, as information seekers awaited calls from patient coordinators about their surgery time. On the day of surgery, questions focused on discharge procedures and post-surgery medication and dosage. After the surgery, participants were concerned about pain management and sought guidance on dos-and-don’ts during recovery (e.g., “washing hair”, “doing yoga”, and “watching TV”).
39 messages from 5 patients and 13 attendants showed signs of anthropomorphizing the bot with addressing the bot as ‘Doctor’, ‘Sir’, or ‘Ma’am’. Many queries were prefaced with greetings (like ‘hi’) and included acknowledgements (like ‘okay’), contributing to the humanization of the bot. In Indic languages, patients employed honorifics when addressing the bot in their messages and during interviews. Notably, two patients with low literacy levels treated the chat as if they were conversing back-and-forth with a human, providing non-actionable status updates, such as P29 telling CataractBot that “The date of the surgery is tomorrow. We will leave home at 8 am from Ayyappa Nagar.”
6.1.4. Multiple Stakeholders
Patients sent a total of 89 messages (4.75.4 message/patient), while attendants sent 254 messages (8.56.8 message/attendant), significantly more than patients, with t(47)=2.04, p¡0.05 (Figure 4C). Among the 309 verified medical and logistical answers, 62.5% were verified by the expert doctor/coordinator and the rest 37.5% by the escalation expert. This highlights the significance of the bot supporting multiple information seeker and information verifier roles. During our interviews, we identified a few reasons for the patients to minimally use the bot, particularly related to demographics. Patients were typically older with lower levels of literacy and tech-literacy compared to their attendants These patients placed trust in the hospital to provide them with all the required information and did not actively seek information themselves. For instance:
“Dad didn’t try… I was the one thinking about it [the surgery]. He was like, “It’s OK! Not everyone needs to bother in the family… (to know) about what will be this, what will be that”. – A3.
Additionally, patients were advised to minimize screen use post-surgery, contributing to their reduced bot interaction. Interestingly, patients sometimes offloaded the knowledge-seeking ‘work’ to their active attendants—for instance, “I have been using it. I told him [the patient] what (answers) I got. So maybe he did not feel the need (to use it).” (A20)—so the involvement of additional stakeholders had a partially negative impact..
Moving on to expert stakeholders, both the doctors and patient coordinators highlighted the significant role of escalation experts. Among the medical questions, 76.2% were verified by operating doctors and 23.7% by the escalation doctor. Doctors mentioned having 2-3 days a week reserved for surgeries, and on these “operating days”, they rarely checked their phones during work hours, leading to more escalations. Despite the crucial role of escalation experts, there was a notable difference in their interactions compared to operating doctors and coordinators. While the latter had previous face-to-face meetings with patients and attendants, there was no such guarantee for escalation experts. As a result, escalation experts lacked knowledge of patients’ context, and hence provided generic responses.
Among the different stakeholders, bot interactions varied. For patients and attendants, the bot operates as a user-driven platform where they initiate conversations by asking questions. In contrast, the bot functions as a bot-driven system for experts, initiating and driving verification-related conversations. Patients and attendants appreciated that CataractBot “does not disturb” them and provides information solely based on the questions asked. However, a few participants suggested the bot should proactively offer relevant information. For example, four attendants wished for regular updates from the chatbot during the surgery while waiting anxiously outside. A9 wanted reminders from the chatbot to ensure patients “take eye drops on time” post-surgery.
6.1.5. Question Time and Response Time
The study was designed for patients and attendants to use the bot seven days pre-surgery and seven days post-surgery. However, due to a bug in the code, the bot was never disabled. Moreover, 3 participants had their surgeries rescheduled, leading them to use the bot for a longer period. Overall, participants sent messages to the bot from 19 days before their surgeries to 24 days post their surgeries (Figure 4B). A majority of messages (71.4%) were sent in the 8-day period from 5 days prior to surgery to 2 days post-surgery with the highest number of messages sent on the day of the surgery (76). 156 messages were sent before the surgery day and 111 post the surgery day. We conducted a one-way ANOVA on the number of messages asked by each patient-attendant group pre-surgery, on the day of surgery, and post-surgery, and found statistical significance with F(2, 60)=3.69, p¡.05. Post-hoc pairwise comparisons with Bonferroni correction indicated a statistically significant difference between ‘pre’ and ‘on’ surgery (t(30)=2.7, p¡.05) and between ‘pre’ and ‘post’ surgery (t(30)=3.2, p¡.05). This suggests that patients and attendants are more inquisitive before surgery, potentially due to heightened anxiety, aligning with prior research (Beisecker and Beisecker, 1990). The higher pre-surgery usage may also be due to the novelty effect. With respect to the types of queries, logistical questions were mainly asked before the surgery, reaching their peak one day before the surgery day (Figure 4B), as anticipated. In contrast, experts expressed that questions on the day of surgery might be unnecessary:
“Multiple questions were asked on the day of surgery, “when he is going to be discharged?” … You are in the hospital… Why to use the bot!… You can walk 10 steps and go to the staff and ask.” – D1.
Regarding the usage patterns of experts, responses were verified throughout the day. The only exception was during days reserved for surgeries. All doctors emphasized a no-phone-use policy during surgeries, stressing the need for minimal messages on those days. A coordinator suggested adding a working-time parameter for CataractBot to restrict message delivery to specified periods. Experts refrained from using their phones in front of patients due to the negative perception associated with using phones for personal reasons. During working hours, the long queue of waiting patients made it challenging for experts to use their phones. These contributed to several queries being escalated. Opinions on using the chatbot after work hours varied: two experts were reluctant, while four were in favor of using it “anytime”.
6.2. Features
6.2.1. Experts-in-the-loop
The novel feature of having human experts verify and update the auto-generated responses was highly appreciated by participants, serving as the key reason for their engagement with the bot. For instance, “That’s why I gave my WhatsApp (number to onboard the bot), because I can talk directly to the doctor.” – A2. Contrary to initial expectations that understanding the icons might be complex, participants found it self-explanatory:
“That’s simple… Even my dad understood. I didn’t tell him the tick mark, question mark… He told me, “okay, doctor verified… Now the tick mark came”… If MY dad was able to use, I think anyone can.” – A13.
However, since participants knew that these messages were being reviewed by doctors and coordinators, four used the bot to send direct messages to them instead of asking questions. For instance, P19 messaged “Dr. Give me a call?”
The length of expert corrections in characters varied significantly (t(91)=5.2, p¡0.05) with 152.4104.9 for medical questions and 50.942.6 for logistical questions. Further analysis revealed that their suggested edits fell into five categories: adding missing information (76.8%), asking clarification question (14.6%), factual corrections (11.0%), asking the patient to visit the hospital (8.8%), and removing extra unnecessary information (3.7%).
Finally, experts did not have access to the final updated answer sent to information seekers after their suggested edits. This design decision limited verification to a single round, minimizing the workload of experts. However, without this transparency, experts were sometimes unsure whether or how CataractBot had understood their corrections, especially when the edits aimed at improving readability and structure rather than solely focusing on accuracy and completeness: “I’d asked to flip [the answer] around… to give information for local anesthesia first then for general anesthesia… Did the bot actually understood and give that as an answer? I donno.” (D3). In the CataractBot system, only the KB update expert had access to the updated answer sent to the information seeker.
6.2.2. Input Modality
Out of the total 343 messages sent by information seekers, 206 (60.1%) were text, 43 (12.5%) audio and 94 (27.4%) related questions. 19 participating patients sent 89 messages (43 text, 26 audio, and 20 taps), while the 30 attendants sent 254 messages (163 text, 17 audio, and 74 taps) (Figure 4C). One-way ANOVAs on message types found statistical significance for attendants (F(2, 87)=13.74, p¡.001), but not for patients (F(2, 54)=0.77, p=0.46). Post-hoc pairwise comparisons, corrected with Holm’s sequential Bonferroni procedure, indicated statistically significant differences between ‘text’ and ‘audio’ (t(29)=4.8, p¡.01) and ‘text’ and ‘taps’ (t(29)=3.0, p¡0.01) for attendants. The minimal usage of audio by attendants could be due to demographic differences, with the majority being under 45 years and well-educated. Audio messages were predominantly used by older semi-literate patients. For instance, “I can’t read (or) write Hindi that well, so asking and listening to audio, I did that.” (P12).
Post-surgery limitations on screen use also made speech input and audio output the preferred modality for patients, with 18 out of 26 audio questions by patients asked post-surgery. 34 participants did not use the audio modality at all, while 31 selected one of the related questions. As P7 stated: “I used the audio message once. After that there was no need for me… I just selected from the options. Click here (shows the phone) and automatically questions will come, and click a question and get the answer.”. While they found the related questions feature convenient and frequently used it, there were instances of dissatisfaction. On eight instances, the answers to suggested questions resulted in IDK responses. This issue arose because the CataractBot system uses an LLM call to suggest three related questions based on the users’ last question, independent of the KB. This approach aimed to uncover gaps in the KB by not restricting questions to the curated KB, and introduce personalization by suggesting questions relevant to the ongoing conversation.
With respect to experts, they could only interact with CataractBot using text. This design choice facilitated a “hygiene check” (D4) for thoroughness and precision in expert’s suggested edits, aligning with previous findings (Li et al., 2023).
6.2.3. Multilinguality
Experts interacted with CataractBot only in English. Among our patients and attendants, the majority (29 out of 49) selected English, while 7 opted for Hindi, 6 Kannada, 5 Tamil, and 2 Telugu. None of the participants changed their initially selected language. Given the challenges of typing in Indic languages on smartphones (Joshi et al., 2011), our non-English participants predominantly relied on speech input and selecting from related questions. In total, patients who chose English typed 35 questions, asked 0 questions in speech, and tapped 6 related questions. In contrast, patients opting for Indic languages typed 8 questions, sent 26 audio questions and tapped 14 related questions in total. A similar trend was observed among attendants. Patients found value in local language interaction and speech as input, helping them use the bot independently: “Without even asking me for help, he (patient) was able to use that bot as it supports Tamil… I felt it’s very useful. Because now, for elders… it is very hard to use apps and things.” – A13.
The audio queries in Indic languages were transcribed and translated to English using Azure AI services. However, this is not highly accurate due to the variability in dialect across India. It is a known issue and prior work (Jain et al., 2018) relied on human wizards to correct such errors. For example, a query by P29 (Kannada) was transcribed and translated to “I got kampu?”. D3, upon listening to the audio query, realized that the patient was asking about a blood clot (as ‘Kempu’ in Kannada means ‘redness’) and was able to respond appropriately. Experts faced challenges in “deciphering” such messages since they only received the English transcription. They suggested providing the original audio recordings of patient queries. D4 added that this would also help extract non-verbal cues, enabling them to respond more appropriately.
It is common for people to type Indic languages in Latin script (instead of using, e.g., Devanagari script for Hindi), a feature not supported by CataractBot due to current translation technology limitations. We found three participants did that. For one of them, the bot responded accurately as the message contained a code-mix of Hindi and English:
A8: agar opration k baad pain ho raha to kya karna hai?
Bot: If the patient is experiencing pain after the surgery, they should report to the doctor immediately…
For the other two, the bot responded with “You seem to be asking the question in a language different from the selected language…”. Additionally, even in the selected language, participants with low literacy and language proficiency struggled to formulate questions (similar to findings in prior research (Jain et al., 2018)). For example, “Tomorrow operation When should come No phone call received yet?” (P18). Such questions were common with 65 questions, of which 26 were asked using audio in an Indic language. Further analysis revealed that the transcription engine was mainly responsible for these errors (19 out of 26), while the remaining 7 were translation errors. Despite these challenges, the robustness of the GPT-4 model enabled the chatbot to answer most (45) of such ill-formulated questions appropriately.
6.2.4. KB Updation
Out of the total 91 responses edited by experts, 72 (79.1%) received approval from the KB update expert (D4) to be added to the KB. For 49.3% of these, the final response generated by the bot (by merging the initial response and expert edit suggestion) underwent additional editing by the KB update expert. These changes were minimal with an average relative Hamming distance of 28.2%, and mainly addressed aspects of “tone and structure”, such as preferring active over passive voice and making responses more generally applicable across patients. This hints that the bot effectively incorporated the expert edit suggestions into the final response.
Experts noticed improvements in the completeness and accuracy of CataractBot’s generated answers over time, thus reducting their workload: “With time, I think the answers did get better… I am mostly just saying yes.” (D3). However, there was an instance where the KB update expert approved incorrect information (i.e., the hospital’s phone number with an extra digit) to be added to the KB. CataractBot provided this to three seekers, before it was fixed. Despite expanding the KB, we identified our bot’s limited knowledge in two medical areas—lens types and their comparisons, and lab tests—and one logistical area—insurance tie-ups. Addressing these required manually adding new documents. D4 suggested the need for a simple way for experts to upload such documents to further expand the KB.
6.3. Privacy and Trust
6.3.1. Personalization vs Privacy
Three participants appreciated that their medical records were not accessible to bot.
“Using personal health information [by the bot] is… a violation of patient’s rights. If their data is added… external agencies can extract that information and might use it for advertising.” – A17.
Three educated participants (holding Master’s degrees) considered personalization as “an unnecessary add-on”. They raised concerns about potential complications associated with sharing PII with LLMs, both in terms of legislation and “privacy threats”. A1, e.g., remarked “I would definitely not want my mom’s data to be out there, all over the internet.”. Moreover, lack of personalization enabled users to use the chatbot for others undergoing cataract surgery: “Two people I know are going to have their cataract surgery soon… I’ll use this bot for them.” (A11).
On the other hand, fueled by the recent hype around AI and LLMs, eight participants expected “precise” and “personalized” answers to their queries with little concern for privacy, and were slightly disappointed with generic responses. This aligns with prior findings indicating minimal privacy concerns related to digital data in the global south, specifically in India (Okolo et al., 2021). Interestingly, three participants did not even consider their medical data to be “private”, as they stated, “It’s okay, it’s nothing personal, to be hidden or something. All have common health issues.” (A14). Moreover, two participants questioned why others would be interested in their medical data, as “it is not useful for anyone else” (P18).
Two educated participants were willing to share their personal medical data if the benefits outweighed the costs. Additionally, two participants wanted personalization, along with the sharing of patient medical data, to be an “opt-in” feature where individuals conduct a cost-benefit analysis and make an informed choice based on their preferences. They believed that a personalized CataractBot would offer more relevant responses and minimize IDK responses.
“If you don’t want to share your information [with the bot], it should say “If you ask personal questions, I won’t have access, so you may have to visit the hospital.”… Somewhere we have to compromise.” – A21.
Five experts were in favor of integrating the chatbot with patient data, including schedules and consultation history, to offer personalized responses. Such integration would not only reduce their workload, but also enhance user trust. D3 suggested collecting basic patient medical history to facilitate customized responses from the bot and doctors: “We don’t even know whether they’re diabetic, they’re hypertensive. So, when they ask “What should I eat and come?”… I ain’t sure what to tell.” (D3). Although the onboarding form attempts to capture such information in the ‘Extra Details’ field, none of our participants provided that information due to time constraints. However, D1 and D2 also voiced skepticism towards the LLMs capability to understand complex patient records accurately. “There are a lot of variables involved, including the surgeon, the particular patient, their case, tests… It’s too much information to process.” (D1). In addition, the KB update expert cautioned that highly personalized responses would result in minimal updates to the KB, potentially increasing CataractBot-related workload for experts and negatively impacting chatbot adoption among experts.
6.3.2. Reliability and Trust
Six information seekers explicitly mentioned trusting the bot because of their trust in the hospital, where they had chosen to undergo surgery. As CataractBot was deployed through Sankara Eye Hospital’s verified Facebook Business account, its default WhatsApp username included the hospital’s name and displayed a green tick indicating ‘verified’ status (Figures 1, 2). Moreover, the integration of the bot with the hospital was reinforced by the patient coordinator advising them to register and use the bot. All these contributed to participants’ trust in the bot.
With respect to bot’s responses, participants gained trust due to the human experts verifying answers (Figure 1F). As A12 stated: “Basically the trust only came when I saw the green tick mark… Before that, I was also in the question mark zone… So I will say it’s a machine… But then somebody is checking it, confirming it… It’s actually then trustworthy.”. The experts-in-the-loop ensured that the bot’s response were more reliable and “not fake” (P18). This is in contrast to a few participants who complained about receiving false information from Google search results.
Although the experts-in-the-loop feature increased trust in final responses, it had a negative impact on perceived bot intelligence. Four participants noted that when an expert corrected CataractBot’s initial answer, trust in the bot’s unverified answers reduced. For instance, “After some time, when the doctor said it is invalid, I was like, okay, should I even trust the bot?” (P19). In total, 29.4% of the bot’s generated answers for medical and logistical queries were marked incorrect by the expert. This could be attributed to the limited custom KB, the experts’ high standards for the tone and structure of answers, or our prompting strategy favoring caution. A9 pointed: “It said that severe cough is dangerous on the surgery day as it increases eye pressure. But the doctor said it is completely fine… [The bot] unnecessarily alarmed me.”
6.3.3. Accountability
CataractBot provides a persistent written record of expert-patient communication, which has implications. Patients mentioned utilizing the information provided by CataractBot as a reference and sharing it with others undergoing cataract surgery. Also, this information can be valuable when discussing their surgery with experts outside of the Sankara Eye Hospital ecosystem: “[Patients] can share the same information with their doctor at their native place… So that [communication with another doctor] becomes a little easier.” (D1). However, it also increased the sense of accountability among experts, making them “somewhat nervous”. They felt the need to be “very lawyer-like in our conversation because we don’t want it to come back and bite us tomorrow.” (D4).
7. Discussion
In this paper, we propose CataractBot, an LLM-powered WhatsApp-based experts-in-the-loop chatbot to address the information needs of patients undergoing cataract surgery. Our user study revealed positive evidence of its usability among the different stakeholders. Patients and attendants appreciated CataractBot for its ability to answer questions anytime with verified responses, which helped establish trust. Additionally, it streamlined communication, saving time by reducing the need for hospital calls/visits. Individuals with low literacy and low-tech literacy found it easy to use because of the multilingual and multimodal support. Experts commended CataractBot for acting as a facilitator, creating a layer between experts and patients. The asynchronous communication feature provided them the flexibility to verify responses at their convenience. We identify critical insights, open questions, and design implications that warrant further attention for such experts-in-the-loop chatbots to succeed on a larger scale. We discuss these aspects below.
7.1. Chatbots in Healthcare and Beyond
7.1.1. Experts-in-the-loop
Several studies have raised concerns about the application of LLMs in healthcare scenarios, due to their inconsistency, potential errors, obscured reasoning, and bias (Denecke et al., 2024; Au Yeung et al., 2023; Fogel and Kvedar, 2018; Bérubé et al., 2021; Iliashenko et al., 2019; Kocaballi et al., 2020). Moreover, users who are less (technologically) literate face a heightened risk of harm (Kapania et al., 2022). To address these concerns and comply with OpenAI’s usage policies (OpenAI, 2024), which state that “tailored medical/health advice cannot be provided without review by a qualified professional”, CataractBot incorporates experts-in-the-loop. This feature ensured reliability and fostered trust, as users knew their operating doctor was verifying the answers. However, instances where users received information from the bot that was subsequently marked as incorrect or incomplete by the expert, their trust in the initial automated answers was reduced. To address this issue, one design approach could be to eliminate the initial automated response to the patient and only send expert-verified answers. Although this might result in long wait times, which could discourage usage of the bot and prompt patients to seek information elsewhere, it is worth exploring. In the future, it would be valuable to compare these two approaches: a hybrid method of synchronous automated answers with asynchronous verification versus a fully asynchronous system with solely verified answers.
7.1.2. Conversational Design
Prior research indicates that synchronous telehealth systems can make patients feel rushed and unable to adequately express their concerns (Gordon et al., 2020; Bhat et al., 2021). In contrast, CataractBot introduces a crucial buffer in the conversation between patients and doctors, allowing them to use the system at their convenience. However, this asynchronicity also posed challenges for both parties. Patients were unsure exactly when they would receive a verified response, while experts received no further feedback from patients after providing a verification. To address this, first, patients and attendants should be made aware of the experts’ availability and informed about the expected delay in verification, particularly when questions are sent outside of working hours, on surgery days, or when experts are on leave. To achieve this, experts should be able to specify their availability in natural language on CataractBot. This will also prevent disturbing experts during their personal time. Second, to provide experts with an acknowledgement after an answer is verified, patients and attendants should be offered a poll: ‘Yes, my question was answered’—which would notify the expert, and ‘No, please give me a call’—which would alert the patient coordinator. This should be approached with caution, as prior work (Weisz et al., 2019; Yeh et al., 2022) suggests that users tend to bypass the automated agent in favour of directly connecting with a human agent via voice call.
7.1.3. Knowledge Base (KB)
For CataractBot, we curated a custom KB. Even within our short deployment, we noticed the KB to be inadequate. For longitudinal deployments, we recommend implementing a simple method for experts to add new documents directly to the KB. Recognizing the impracticality of manual updates (after approval by the KB update expert) in rapidly changing environments, we also propose automatic updates based on the consensus among doctors in the edited answers. Lastly, as the KB updates, a method is needed to evaluate CataractBot’s answers over time. Tan et al. (2024) found that GPT-4’s evaluations of AI-generated responses align closely with human clinician assessments, demonstrating the LLM’s potential not only for generating but also for validating answers. Incorporating an LLM as an evaluator has the potential to further improve the process of automatically expanding the KB while ensuring that the quality of the responses remains intact.
7.1.4. Beyond Cataract Surgery
Cataract surgery, with its brief duration, straightforward procedure, high success rate, and extensive documentation, provided an ideal low-stakes environment for building and testing our experts-in-the-loop chatbot. A similar chatbot could be even more valuable for information seekers undergoing complex, atypical, or long-term treatments, such as glaucoma surgery, cancer treatment, or childbirth. With our open-sourced framework for developing experts-in-the-loop bots, this should be feasible. Furthermore, the applicability of such a system need not be confined to the medical domain. Other fields, such as legal, finance, education, or agriculture, could benefit from similar chatbots. In these scenarios, end-users (e.g., defendants, taxpayers, students, or farmers) could receive synchronous responses from an LLM operating on a custom KB and subsequently obtain verified responses asynchronously from experts (e.g., lawyers, accountants, teachers/career counselors, or agricultural experts, respectively). Expert-mediated bots can even be used to upskill less-trained experts, such as patient coordinators or paralegals. For instance, a maternal and child health bot could serve as a valuable learning resource for community health workers in rural or semi-urban areas, who typically have minimal (medical) training. Such a bot can help them provide accurate information to patients and their families while also boosting their confidence and competence.
7.2. Design Implications
7.2.1. Personalization
Personalized responses present a double-edged sword—while they can enhance the adoption and utility of CataractBot by providing more relevant information, they may also raise privacy concerns among certain users, potentially reducing usage. Additionally, personalized responses could increase the workload of experts, as they would need to refer to specific patient documents for verification. Many of these verification edits by experts may not be suitable for inclusion in the KB due to their patient-specific nature. Given our demographic of low literate and low tech-literate users, they might require assistance in making informed decisions about consenting to the use of their health data (Okolo et al., 2021). Also, adopting a binary approach–either sharing all medical records or none–may not align with user preferences. Instead, users should have the agency to selectively share records they deem relevant and acceptable for the bot’s access, in line with the philosophy of the National Digital Health Mission (NDHM) (National Health Authority and Family Welfare, 2020) in India. A solution could be to ask users to manually upload relevant documents (such as consultation sheets, discharge summaries, and test reports) or textually enter details they wish the bot to access. This approach can empower users to control the information available to the bot for a more personalized experience. Currently, CataractBot operates as a user-driven system. With personalization features, we propose expanding its role to drive interactions at crucial points during the patient’s journey, e.g., providing medication reminders.
7.2.2. Reduce Workload of Experts
Deploying CataractBot in the hospital increased the workload of patient coordinators (due to onboarding and verification) and doctors (due to verification and KB updates). To alleviate their burden, we propose ways to streamline CataractBot-related tasks and to use CataractBot to support them in their day-to-day activities. First, to simplify patient onboarding, QR codes can be installed at key locations in the hospital. Users can scan the code, self-onboard by providing necessary information (such as language preference, operating doctor, and surgery date), and instantly engage with the bot. Users should receive a brief tutorial video on interaction guidelines, emphasizing that the bot is not intended for emergencies. Second, CataractBot should be utilized to deliver general information about the surgery and provide status updates to patients and attendants, thereby assisting experts in managing their other responsibilities. This would enable experts focus on more specific and meaningful conversations with their patients during their limited face-to-face time. Third, to reduce the verification load for experts, we propose a system to pre-verify frequently asked questions automatically based on experts’ repeated confirmations of correctness. Finally, regarding audio messages, the lack of non-verbal cues hindered experts’ ability to deliver care (similar to (Li et al., 2023)). To address the limitations of translation and transcription technologies, the information seeker’s original audio clip should also be shared with the experts.
7.2.3. Usefulness
Attfield et al. (2006) reported that patients seek information to prepare for a doctor consultation, verify a diagnosis or treatment plan, and assess the need for clinical intervention. However, access to CataractBot is currently provided only post-scheduling of the surgery. Introducing the bot before this might conflict with hospital policies, as patients could potentially cancel surgeries after using the bot, resulting in resource wastage for both the bot and experts. During bot interactions, a few participants expressed frustration with IDK responses (similar to previous findings (Xiao et al., 2023)), leading to discontinuation of usage. To address this, we suggest enhancing IDK responses by providing general information about the topic, offering tentative recommendations with disclaimers, and guiding users to alternative information sources. Lastly, considering the visual capabilities of recent LLM models like GPT-4V (a multimodal model), CataractBot should be enhanced to handle image-based inputs, aligning with recommendations from previous work (Jain et al., 2018). For instance, patients could upload images of medications instead of providing written details. While this could improve the user experience for information seekers, it may also increase the workload for experts.
7.3. Limitations
We acknowledge several limitations of this work. First, the results should be interpreted considering the specificities of the study, including the relatively small sample size, users’ first encounter with a medical chatbot, and the cataract treatment scenario. The positive user responses and other conclusions should be validated by future, preferably longitudinal studies. Second, system-level analyses, such as scalability testing and power measurement, are needed before scaling the proposed solution to the masses. Third, CataractBot uses expert responses to incrementally update the custom KB; however, the new information may conflict with the bot’s older KB. This issue requires further investigation. Finally, we note the potential risk of disseminating inaccurate medical information at scale with chatbot and LLM technologies. Therefore, stronger methods are needed to develop, review, and update the KB, as well as to regularly evaluate the accuracy of bot’s responses.
8. Conclusion
Building a system that addresses patients’ information needs by providing expert-verified information is an open problem in healthcare. To tackle this, we propose a novel solution—a LLM-powered experts-in-the-loop chatbot framework, that utilizes RAG over a custom KB to provide synchronous responses, and utilizes experts to provide verified responses asynchronously. Our in-the-wild study involving 49 participants (patients and attendants) and 6 experts (4 doctors and 2 coordinators) demonstrated the positive impact of this technological intervention. Patients not only trusted the chatbot’s response, but were also willing to wait, understanding that their busy doctors were verifying them. Simultaneously, doctors and coordinators, despite their hectic schedules, made time during breaks and at home to help patients by verifying LLM-generated responses. The favorable feedback from various stakeholders in the healthcare ecosystem indicates that a chatbot, delivered through ubiquitous smartphones and WhatsApp, can effectively enhance information access in critical healthcare settings, even for individuals with limited literacy and technology experience.
Acknowledgements.
Many thanks to all the participants for their time and patience.References
- (1)
- ait ([n. d.]) [n. d.]. An AI service for real-time document and text translation. https://www.microsoft.com/en-us/ai/ai-translator Accessed: 2024-02-01.
- Abd-Alrazaq et al. (2020) Alaa Abd-Alrazaq, Zeineb Safi, Mohannad Alajlani, Jim Warren, Mowafa Househ, and Kerstin Denecke. 2020. Technical metrics used to evaluate health care chatbots: scoping review. Journal of medical Internet research 22, 6 (2020), e18301.
- Ahuja et al. (2023) Kabir Ahuja, Harshita Diddee, Rishav Hada, Millicent Ochieng, Krithika Ramesh, Prachi Jain, Akshay Nambi, Tanuja Ganu, Sameer Segal, Maxamed Axmed, Kalika Bali, and Sunayana Sitaram. 2023. MEGA: Multilingual Evaluation of Generative AI. arXiv:2303.12528 [cs.CL]
- Akerkar and Bichile (2004) Shashank M Akerkar and Lata S Bichile. 2004. Doctor patient relationship: changing dynamics in the information age. Journal of postgraduate medicine 50, 2 (2004), 120–122.
- Arora et al. (2008) Neeraj K Arora, Bradford W Hesse, Barbara K Rimer, Kasisomayajula Viswanath, Marla L Clayman, and Robert T Croyle. 2008. Frustrated and confused: the American public rates its cancer-related information-seeking experiences. Journal of general internal medicine 23 (2008), 223–228.
- Attfield et al. (2006) Simon J. Attfield, Anne Adams, and Ann Blandford. 2006. Patient information needs: pre- and post-consultation. Health Informatics Journal 12, 2 (2006), 165–177. https://doi.org/10.1177/1460458206063811 arXiv:https://doi.org/10.1177/1460458206063811 PMID: 17023406.
- Au Yeung et al. (2023) Joshua Au Yeung, Zeljko Kraljevic, Akish Luintel, Alfred Balston, Esther Idowu, Richard J Dobson, and James T Teo. 2023. AI chatbots not yet ready for clinical use. Frontiers in Digital Health 5 (2023), 60.
- Beasley (2009) Robert E Beasley. 2009. Short message service (SMS) texting symbols: A functional analysis of 10,000 cellular phone text messages. The Reading Matrix 9, 2 (2009).
- Beisecker and Beisecker (1990) Analee E. Beisecker and Thomas D. Beisecker. 1990. Patient Information-Seeking Behaviors When Communicating with Doctors. Medical Care 28, 1 (1990), 19–28. http://www.jstor.org/stable/3765618
- Bernstein et al. (2023) Isaac A Bernstein, Youchen Victor Zhang, Devendra Govil, Iyad Majid, Robert T Chang, Yang Sun, Ann Shue, Jonathan C Chou, Emily Schehlein, Karen L Christopher, et al. 2023. Comparison of ophthalmologist and large language model chatbot responses to online patient eye care questions. JAMA Network Open 6, 8 (2023), e2330320–e2330320.
- Bérubé et al. (2021) Caterina Bérubé, Zsolt Ferenc Kovacs, Elgar Fleisch, and Tobias Kowatsch. 2021. Reliability of Commercial Voice Assistants’ Responses to Health-Related Questions in Noncommunicable Disease Management: Factorial Experiment Assessing Response Rate and Source of Information. Journal of Medical Internet Research 23, 12 (2021), e32161.
- Bhat et al. (2021) Karthik S Bhat, Mohit Jain, and Neha Kumar. 2021. Infrastructuring Telehealth in (In)Formal Patient-Doctor Contexts. Proc. ACM Hum.-Comput. Interact. 5, CSCW2, Article 323 (oct 2021), 28 pages. https://doi.org/10.1145/3476064
- Bickmore et al. (2009) Timothy W Bickmore, Laura M Pfeifer, and Brian W Jack. 2009. Taking the time to care: empowering low health literacy hospital patients with virtual nurse agents. In Proceedings of the SIGCHI conference on human factors in computing systems. 1265–1274.
- Bodenheimer and Smith (2013) Thomas S Bodenheimer and Mark D Smith. 2013. Primary care: proposed solutions to the physician shortage without training more physicians. Health Affairs 32, 11 (2013), 1881–1886.
- Borkar (2023) Jaydeep Borkar. 2023. What can we learn from Data Leakage and Unlearning for Law? arXiv preprint arXiv:2307.10476 (2023).
- Byrne and Heyman (1997) G Byrne and R Heyman. 1997. Patient anxiety in the accident and emergency department. J. Clin. Nurs. 6, 4 (July 1997), 289–295.
- Cajander and Grünloh (2019a) Åsa Cajander and Christiane Grünloh. 2019a. Electronic Health Records Are More Than a Work Tool: Conflicting Needs of Direct and Indirect Stakeholders. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (Glasgow, Scotland Uk) (CHI ’19). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3290605.3300865
- Cajander and Grünloh (2019b) Åsa Cajander and Christiane Grünloh. 2019b. Electronic Health Records Are More Than a Work Tool: Conflicting Needs of Direct and Indirect Stakeholders. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (Glasgow, Scotland Uk) (CHI ’19). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3290605.3300865
- Chandwani and Kulkarni (2016) Rajesh Chandwani and Vaibhavi Kulkarni. 2016. Who’s the Doctor? Physicians’ Perception of Internet Informed Patients in India. In Proceedings of the 2016 CHI Conference on Human Factors in Computing Systems (San Jose, California, USA) (CHI ’16). Association for Computing Machinery, New York, NY, USA, 3091–3102. https://doi.org/10.1145/2858036.2858500
- Christalle et al. (2019) Eva Christalle, Jördis Zill, Wiebke Frerichs, Martin Härter, Yvonne Nestoriuc, Jörg Dirmaier, and Isabelle Scholl. 2019. Assessment of patient information needs: A systematic review of measures. PLOS ONE 14 (01 2019), e0209165. https://doi.org/10.1371/journal.pone.0209165
- Chung (2017) Chia-Fang Chung. 2017. Supporting patient-provider communication and engagement with personal informatics data. In Proceedings of the 2017 ACM International Joint Conference on Pervasive and Ubiquitous Computing and Proceedings of the 2017 ACM International Symposium on Wearable Computers (Maui, Hawaii) (UbiComp ’17). Association for Computing Machinery, New York, NY, USA, 335–338. https://doi.org/10.1145/3123024.3123197
- Civan-Hartzler et al. (2010) Andrea Civan-Hartzler, David W. McDonald, Chris Powell, Meredith M. Skeels, Marlee Mukai, and Wanda Pratt. 2010. Bringing the field into focus: user-centered design of a patient expertise locator. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Atlanta, Georgia, USA) (CHI ’10). Association for Computing Machinery, New York, NY, USA, 1675–1684. https://doi.org/10.1145/1753326.1753577
- Clarke et al. (2016a) Martina A Clarke, Joi L Moore, Linsey M Steege, Richelle J Koopman, Jeffery L Belden, Shannon M Canfield, Susan E Meadows, Susan G Elliott, and Min Soon Kim. 2016a. Health information needs, sources, and barriers of primary care patients to achieve patient-centered care: A literature review. Health Informatics Journal 22, 4 (2016), 992–1016. https://doi.org/10.1177/1460458215602939 arXiv:https://doi.org/10.1177/1460458215602939 PMID: 26377952.
- Clarke et al. (2016b) Martina A Clarke, Joi L Moore, Linsey M Steege, Richelle J Koopman, Jeffery L Belden, Shannon M Canfield, Susan E Meadows, Susan G Elliott, and Min Soon Kim. 2016b. Health information needs, sources, and barriers of primary care patients to achieve patient-centered care: A literature review. Health Informatics Journal 22, 4 (2016), 992–1016. https://doi.org/10.1177/1460458215602939 arXiv:https://doi.org/10.1177/1460458215602939 PMID: 26377952.
- Dahlén and Janson (2002) Inger Dahlén and Christer Janson. 2002. Anxiety and depression are related to the outcome of emergency treatment in patients with obstructive pulmonary disease. Chest 122, 5 (Nov. 2002), 1633–1637.
- De Choudhury et al. (2014) Munmun De Choudhury, Meredith Ringel Morris, and Ryen W. White. 2014. Seeking and sharing health information online: comparing search engines and social media. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Toronto, Ontario, Canada) (CHI ’14). Association for Computing Machinery, New York, NY, USA, 1365–1376. https://doi.org/10.1145/2556288.2557214
- Denecke et al. (2024) Kerstin Denecke, Richard May, and Octavio Rivera Romero. 2024. Potential of Large Language Models in Health Care: Delphi Study. J Med Internet Res 26 (13 May 2024), e52399. https://doi.org/10.2196/52399
- Desai and Chin (2023) Smit Desai and Jessie Chin. 2023. OK Google, Let’s Learn: Using Voice User Interfaces for Informal Self-Regulated Learning of Health Topics among Younger and Older Adults. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿Hamburg¡/city¿, ¡country¿Germany¡/country¿, ¡/conf-loc¿) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 847, 21 pages. https://doi.org/10.1145/3544548.3581507
- Ding et al. (2020) Xianghua Ding, Xinning Gui, Xiaojuan Ma, Zhaofei Ding, and Yunan Chen. 2020. Getting the Healthcare We Want: The Use of Online ”Ask the Doctor” Platforms in Practice. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3313831.3376699
- Duffy et al. (2004) F Duffy, Geoffrey Gordon, Gerald Whelan, Kathy Cole-Kelly, Richard Frankel, Natalie Buffone, Stephanie Lofton, MaryAnne Wallace, Leslie Goode, and Lynn Langdon. 2004. Assessing Competence in Communication and Interpersonal Skills: The Kalamazoo II Report. Academic medicine : journal of the Association of American Medical Colleges 79 (07 2004), 495–507. https://doi.org/10.1097/00001888-200406000-00002
- Eriksson-Liebon et al. (2021) Magda Eriksson-Liebon, Susanne Roos, and Ingrid Hellström. 2021. Patients’ expectations and experiences of being involved in their own care in the emergency department: A qualitative interview study. Journal of clinical nursing 30, 13-14 (2021), 1942–1952.
- Fadhil (2018) Ahmed Fadhil. 2018. A conversational interface to improve medication adherence: towards AI support in patient’s treatment. arXiv preprint arXiv:1803.09844 (2018).
- Ferré et al. (2020) Fabrice Ferré, Nicolas Boeschlin, Bruno Bastiani, Adeline Castel, Anne Ferrier, Laetitia Bosch, Fabrice Muscari, Matt Kurrek, Olivier Fourcade, Antoine Piau, and Vincent Minville. 2020. Improving Provision of Preanesthetic Information Through Use of the Digital Conversational Agent “MyAnesth”: Prospective Observational Trial. J Med Internet Res 22, 12 (4 Dec 2020), e20455. https://doi.org/10.2196/20455
- Findlater et al. (2009) Leah Findlater, Ravin Balakrishnan, and Kentaro Toyama. 2009. Comparing Semiliterate and Illiterate Users’ Ability to Transition from Audio+text to Text-Only Interaction. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Boston, MA, USA) (CHI ’09). Association for Computing Machinery, New York, NY, USA, 1751–1760. https://doi.org/10.1145/1518701.1518971
- Fochsen et al. (2006) Grethe Fochsen, Kirti Deshpande, and Anna Thorson. 2006. Power Imbalance and Consumerism in the Doctor-Patient Relationship: Health Care Providers’ Experiences of Patient Encounters in a Rural District in India. Qualitative Health Research 16, 9 (2006), 1236–1251. https://doi.org/10.1177/1049732306293776 arXiv:https://doi.org/10.1177/1049732306293776 PMID: 17038755.
- Fogel and Kvedar (2018) Alexander L Fogel and Joseph C Kvedar. 2018. Artificial intelligence powers digital medicine. NPJ digital medicine 1, 1 (2018), 5.
- Gordon et al. (2020) Howard S. Gordon, Pooja Solanki, Barbara G. Bokhour, and Ravi K. Gopal. 2020. “I’m Not Feeling Like I’m Part of the Conversation” Patients’ Perspectives on Communicating in Clinical Video Telehealth Visits. Journal of General Internal Medicine 35, 6 (01 Jun 2020), 1751–1758. https://doi.org/10.1007/s11606-020-05673-w
- Grönvall and Verdezoto (2013) Erik Grönvall and Nervo Verdezoto. 2013. Beyond self-monitoring: understanding non-functional aspects of home-based healthcare technology. In Proceedings of the 2013 ACM International Joint Conference on Pervasive and Ubiquitous Computing (Zurich, Switzerland) (UbiComp ’13). Association for Computing Machinery, New York, NY, USA, 587–596. https://doi.org/10.1145/2493432.2493495
- Henry et al. (2022) Katharine E Henry, Rachel Kornfield, Anirudh Sridharan, Robert C Linton, Catherine Groh, Tony Wang, Albert Wu, Bilge Mutlu, and Suchi Saria. 2022. Human–machine teaming is key to AI adoption: clinicians’ experiences with a deployed machine learning system. NPJ digital medicine 5, 1 (2022), 97.
- Holmes et al. (2023) Samuel Holmes, Raymond Bond, Anne Moorhead, Jane Zheng, Vivien Coates, and Michael McTear. 2023. Towards Validating a Chatbot Usability Scale. In Design, User Experience, and Usability, Aaron Marcus, Elizabeth Rosenzweig, and Marcelo M. Soares (Eds.). Springer Nature Switzerland, Cham, 321–339.
- Iliashenko et al. (2019) Oksana Iliashenko, Zilia Bikkulova, and Alissa Dubgorn. 2019. Opportunities and challenges of artificial intelligence in healthcare. In E3S Web of Conferences, Vol. 110. EDP Sciences, 02028.
- Ittarat et al. (2023) Mantapond Ittarat, Wisit Cheungpasitporn, and Sunee Chansangpetch. 2023. Personalized Care in Eye Health: Exploring Opportunities, Challenges, and the Road Ahead for Chatbots. Journal of Personalized Medicine 13, 12 (2023), 1679.
- Jain et al. (2018) Mohit Jain, Pratyush Kumar, Ishita Bhansali, Q. Vera Liao, Khai Truong, and Shwetak Patel. 2018. FarmChat: A Conversational Agent to Answer Farmer Queries. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 2, 4, Article 170 (dec 2018), 22 pages. https://doi.org/10.1145/3287048
- Jo et al. (2023) Eunkyung Jo, Daniel A Epstein, Hyunhoon Jung, and Young-Ho Kim. 2023. Understanding the benefits and challenges of deploying conversational AI leveraging large language models for public health intervention. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–16.
- Johansson et al. (2020a) Anders Johansson, Monica Larsson, and Bodil Ivarsson. 2020a. General practitioners’ experiences of digital written patient dialogues: a pilot study using a mixed method. Journal of Primary Care & Community Health 11 (2020), 2150132720909656.
- Johansson et al. (2020b) Anders Johansson, Monica Larsson, and Bodil Ivarsson. 2020b. Patients’ experiences with a digital primary health care concept using written dialogues: a pilot study. Journal of Primary Care & Community Health 11 (2020), 2150132720910564.
- Joseph-Williams et al. (2014) Natalie Joseph-Williams, Glyn Elwyn, and Adrian Edwards. 2014. Knowledge is not power for patients: A systematic review and thematic synthesis of patient-reported barriers and facilitators to shared decision making. Patient Education and Counseling 94, 3 (2014), 291–309. https://doi.org/10.1016/j.pec.2013.10.031
- Joshi et al. (2011) Anirudha Joshi, Girish Dalvi, Manjiri Joshi, Prasad Rashinkar, and Aniket Sarangdhar. 2011. Design and evaluation of Devanagari virtual keyboards for touch screen mobile phones. In Proceedings of the 13th International Conference on Human Computer Interaction with Mobile Devices and Services. 323–332.
- Kapania et al. (2022) Shivani Kapania, Oliver Siy, Gabe Clapper, Azhagu Meena SP, and Nithya Sambasivan. 2022. ”Because AI is 100% right and safe”: User Attitudes and Sources of AI Authority in India. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿New Orleans¡/city¿, ¡state¿LA¡/state¿, ¡country¿USA¡/country¿, ¡/conf-loc¿) (CHI ’22). Association for Computing Machinery, New York, NY, USA, Article 158, 18 pages. https://doi.org/10.1145/3491102.3517533
- Kinnersley et al. (2007) P Kinnersley, A Edwards, K Hood, N Cadbury, R Ryan, H Prout, D Owen, F Macbeth, P Butow, and C Butler. 2007. Interventions before consultations for helping patients address their information needs. Cochrane Database Syst. Rev. 2010, 3 (July 2007), CD004565.
- Klasnja et al. (2010a) Predrag Klasnja, Andrea Civan Hartzler, Kent T. Unruh, and Wanda Pratt. 2010a. Blowing in the Wind: Unanchored Patient Information Work during Cancer Care. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Atlanta, Georgia, USA) (CHI ’10). Association for Computing Machinery, New York, NY, USA, 193–202. https://doi.org/10.1145/1753326.1753355
- Klasnja et al. (2010b) Predrag Klasnja, Andrea Civan Hartzler, Kent T. Unruh, and Wanda Pratt. 2010b. Blowing in the wind: unanchored patient information work during cancer care. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Atlanta, Georgia, USA) (CHI ’10). Association for Computing Machinery, New York, NY, USA, 193–202. https://doi.org/10.1145/1753326.1753355
- Kocaballi et al. (2020) A Baki Kocaballi, Kiran Ijaz, Liliana Laranjo, Juan C Quiroz, Dana Rezazadegan, Huong Ly Tong, Simon Willcock, Shlomo Berkovsky, and Enrico Coiera. 2020. Envisioning an artificial intelligence documentation assistant for future primary care consultations: A co-design study with general practitioners. Journal of the American Medical Informatics Association 27, 11 (08 2020), 1695–1704. https://doi.org/10.1093/jamia/ocaa131 arXiv:https://academic.oup.com/jamia/article-pdf/27/11/1695/34363907/ocaa131.pdf
- Kumar et al. (2023) Harsh Kumar, Yiyi Wang, Jiakai Shi, Ilya Musabirov, Norman AS Farb, and Joseph Jay Williams. 2023. Exploring the Use of Large Language Models for Improving the Awareness of Mindfulness. In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems. 1–7.
- Kurtz (2002) Suzanne M. Kurtz. 2002. Doctor-Patient Communication: Principles and Practices. Canadian Journal of Neurological Sciences / Journal Canadien des Sciences Neurologiques 29, S2 (2002), S23–S29. https://doi.org/10.1017/S0317167100001906
- Leong et al. (2018) Zeina Atrash Leong, Michael S. Horn, Lisa Thaniel, and Emily Meier. 2018. Inspiring AWE: Transforming Clinic Waiting Rooms into Informal Learning Environments with Active Waiting Education. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (Montreal QC, Canada) (CHI ’18). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3173574.3173672
- Lewis et al. (2020) Patrick Lewis, Ethan Perez, Aleksandra Piktus, Fabio Petroni, Vladimir Karpukhin, Naman Goyal, Heinrich Küttler, Mike Lewis, Wen-tau Yih, Tim Rocktäschel, Sebastian Riedel, and Douwe Kiela. 2020. Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks. In Advances in Neural Information Processing Systems, H. Larochelle, M. Ranzato, R. Hadsell, M.F. Balcan, and H. Lin (Eds.), Vol. 33. Curran Associates, Inc., 9459–9474. https://proceedings.neurips.cc/paper_files/paper/2020/file/6b493230205f780e1bc26945df7481e5-Paper.pdf
- Li et al. (2023) Brenna Li, Tetyana Skoropad, Puneet Seth, Mohit Jain, Khai Truong, and Alex Mariakakis. 2023. Constraints and Workarounds to Support Clinical Consultations in Synchronous Text-based Platforms. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿Hamburg¡/city¿, ¡country¿Germany¡/country¿, ¡/conf-loc¿) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 342, 17 pages. https://doi.org/10.1145/3544548.3581014
- Lotfi et al. (2019) Mojgan Lotfi, Vahid Zamanzadeh, Leila Valizadeh, and Mohammad Khajehgoodari. 2019. Assessment of nurse–patient communication and patient satisfaction from nursing care. Nursing open 6, 3 (2019), 1189–1196.
- Ma et al. (2023) Xiumei Ma, Yunxing Liu, Pengfei Zhang, Rongtao Qi, and Fanbo Meng. 2023. Understanding online health information seeking behavior of older adults: A social cognitive perspective. Frontiers in Public Health 11 (2023), 1147789.
- Makarem and Antoun (2016) Nisrine N. Makarem and Jumana Antoun. 2016. Email communication in a developing country: different family physician and patient perspectives. Libyan Journal of Medicine 11, 1 (2016), 32679. https://doi.org/10.3402/ljm.v11.32679 arXiv:https://doi.org/10.3402/ljm.v11.32679 PMID: 28349842.
- Malki et al. (2024) Lisa Mekioussa Malki, Dilisha Patel, and Aneesha Singh. 2024. ”The Headline Was So Wild That I Had To Check”: An Exploration of Women’s Encounters With Health Misinformation on Social Media. Proc. ACM Hum.-Comput. Interact. 8, CSCW1, Article 128 (apr 2024), 26 pages. https://doi.org/10.1145/3637405
- Mamykina et al. (2008) Lena Mamykina, Elizabeth Mynatt, Patricia Davidson, and Daniel Greenblatt. 2008. MAHI: investigation of social scaffolding for reflective thinking in diabetes management. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Florence, Italy) (CHI ’08). Association for Computing Machinery, New York, NY, USA, 477–486. https://doi.org/10.1145/1357054.1357131
- Moden (2022) Rebecca Moden. 2022. Blank Page Syndrome And How To Beat It. Retrieved Sept 1, 2023 from https://www.ool.co.uk/blog/blank-page-syndrome-and-how-to-beat-it/
- Mols and Pridmore (2021) Anouk Mols and Jason Pridmore. 2021. Always available via WhatsApp: Mapping everyday boundary work practices and privacy negotiations. Mobile Media & Communication 9, 3 (2021), 422–440. https://doi.org/10.1177/2050157920970582
- Montagna et al. (2023) Sara Montagna, Stefano Ferretti, Lorenz Cuno Klopfenstein, Antonio Florio, and Martino Francesco Pengo. 2023. Data decentralisation of LLM-based chatbot systems in chronic disease self-management. In Proceedings of the 2023 ACM Conference on Information Technology for Social Good. 205–212.
- National Health Authority and Family Welfare (2020) Ministry of Health National Health Authority and Government of India Family Welfare. 2020. National Digital Health Mission: Strategy Overview. https://www.niti.gov.in/sites/default/files/2023-02/ndhm_strategy_overview.pdf
- Office of the Registrar General & Census Commissioner (2011) Government of India Office of the Registrar General & Census Commissioner, India; Ministry of Home Affairs. 2011. Census of India 2011: C-16: Population by mother tongue, Karnataka - 2011. https://censusindia.gov.in/nada/index.php/catalog/10208/study-description Accessed: 2024-07-24.
- Okolo et al. (2021) Chinasa T. Okolo, Srujana Kamath, Nicola Dell, and Aditya Vashistha. 2021. “It Cannot Do All of My Work”: Community Health Worker Perceptions of AI-Enabled Mobile Health Applications in Rural India. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (Yokohama, Japan) (CHI ’21). Association for Computing Machinery, New York, NY, USA, Article 701, 20 pages. https://doi.org/10.1145/3411764.3445420
- OpenAI (2024) OpenAI. 2024. Usage Policies. https://openai.com/policies/usage-policies/ Accessed: 2024-06-18.
- OpenAI et al. (2023) OpenAI, Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, Red Avila, Igor Babuschkin, Suchir Balaji, Valerie Balcom, Paul Baltescu, Haiming Bao, Mo Bavarian, Jeff Belgum, Irwan Bello, Jake Berdine, Gabriel Bernadett-Shapiro, Christopher Berner, Lenny Bogdonoff, Oleg Boiko, Madelaine Boyd, Anna-Luisa Brakman, Greg Brockman, Tim Brooks, Miles Brundage, Kevin Button, Trevor Cai, Rosie Campbell, Andrew Cann, Brittany Carey, Chelsea Carlson, Rory Carmichael, Brooke Chan, Che Chang, Fotis Chantzis, Derek Chen, Sully Chen, Ruby Chen, Jason Chen, Mark Chen, Ben Chess, Chester Cho, Casey Chu, Hyung Won Chung, Dave Cummings, Jeremiah Currier, Yunxing Dai, Cory Decareaux, Thomas Degry, Noah Deutsch, Damien Deville, Arka Dhar, David Dohan, Steve Dowling, Sheila Dunning, Adrien Ecoffet, Atty Eleti, Tyna Eloundou, David Farhi, Liam Fedus, Niko Felix, Simón Posada Fishman, Juston Forte, Isabella Fulford, Leo Gao, Elie Georges, Christian Gibson, Vik Goel, Tarun Gogineni, Gabriel Goh, Rapha Gontijo-Lopes, Jonathan Gordon, Morgan Grafstein, Scott Gray, Ryan Greene, Joshua Gross, Shixiang Shane Gu, Yufei Guo, Chris Hallacy, Jesse Han, Jeff Harris, Yuchen He, Mike Heaton, Johannes Heidecke, Chris Hesse, Alan Hickey, Wade Hickey, Peter Hoeschele, Brandon Houghton, Kenny Hsu, Shengli Hu, Xin Hu, Joost Huizinga, Shantanu Jain, Shawn Jain, Joanne Jang, Angela Jiang, Roger Jiang, Haozhun Jin, Denny Jin, Shino Jomoto, Billie Jonn, Heewoo Jun, Tomer Kaftan, Łukasz Kaiser, Ali Kamali, Ingmar Kanitscheider, Nitish Shirish Keskar, Tabarak Khan, Logan Kilpatrick, Jong Wook Kim, Christina Kim, Yongjik Kim, Hendrik Kirchner, Jamie Kiros, Matt Knight, Daniel Kokotajlo, Łukasz Kondraciuk, Andrew Kondrich, Aris Konstantinidis, Kyle Kosic, Gretchen Krueger, Vishal Kuo, Michael Lampe, Ikai Lan, Teddy Lee, Jan Leike, Jade Leung, Daniel Levy, Chak Ming Li, Rachel Lim, Molly Lin, Stephanie Lin, Mateusz Litwin, Theresa Lopez, Ryan Lowe, Patricia Lue, Anna Makanju, Kim Malfacini, Sam Manning, Todor Markov, Yaniv Markovski, Bianca Martin, Katie Mayer, Andrew Mayne, Bob McGrew, Scott Mayer McKinney, Christine McLeavey, Paul McMillan, Jake McNeil, David Medina, Aalok Mehta, Jacob Menick, Luke Metz, Andrey Mishchenko, Pamela Mishkin, Vinnie Monaco, Evan Morikawa, Daniel Mossing, Tong Mu, Mira Murati, Oleg Murk, David Mély, Ashvin Nair, Reiichiro Nakano, Rajeev Nayak, Arvind Neelakantan, Richard Ngo, Hyeonwoo Noh, Long Ouyang, Cullen O’Keefe, Jakub Pachocki, Alex Paino, Joe Palermo, Ashley Pantuliano, Giambattista Parascandolo, Joel Parish, Emy Parparita, Alex Passos, Mikhail Pavlov, Andrew Peng, Adam Perelman, Filipe de Avila Belbute Peres, Michael Petrov, Henrique Ponde de Oliveira Pinto, Michael, Pokorny, Michelle Pokrass, Vitchyr Pong, Tolly Powell, Alethea Power, Boris Power, Elizabeth Proehl, Raul Puri, Alec Radford, Jack Rae, Aditya Ramesh, Cameron Raymond, Francis Real, Kendra Rimbach, Carl Ross, Bob Rotsted, Henri Roussez, Nick Ryder, Mario Saltarelli, Ted Sanders, Shibani Santurkar, Girish Sastry, Heather Schmidt, David Schnurr, John Schulman, Daniel Selsam, Kyla Sheppard, Toki Sherbakov, Jessica Shieh, Sarah Shoker, Pranav Shyam, Szymon Sidor, Eric Sigler, Maddie Simens, Jordan Sitkin, Katarina Slama, Ian Sohl, Benjamin Sokolowsky, Yang Song, Natalie Staudacher, Felipe Petroski Such, Natalie Summers, Ilya Sutskever, Jie Tang, Nikolas Tezak, Madeleine Thompson, Phil Tillet, Amin Tootoonchian, Elizabeth Tseng, Preston Tuggle, Nick Turley, Jerry Tworek, Juan Felipe Cerón Uribe, Andrea Vallone, Arun Vijayvergiya, Chelsea Voss, Carroll Wainwright, Justin Jay Wang, Alvin Wang, Ben Wang, Jonathan Ward, Jason Wei, CJ Weinmann, Akila Welihinda, Peter Welinder, Jiayi Weng, Lilian Weng, Matt Wiethoff, Dave Willner, Clemens Winter, Samuel Wolrich, Hannah Wong, Lauren Workman, Sherwin Wu, Jeff Wu, Michael Wu, Kai Xiao, Tao Xu, Sarah Yoo, Kevin Yu, Qiming Yuan, Wojciech Zaremba, Rowan Zellers, Chong Zhang, Marvin Zhang, Shengjia Zhao, Tianhao Zheng, Juntang Zhuang, William Zhuk, and Barret Zoph. 2023. GPT-4 Technical Report. arXiv:arXiv:2303.08774
- Osman Andersen et al. (2021) Tariq Osman Andersen, Francisco Nunes, Lauren Wilcox, Elizabeth Kaziunas, Stina Matthiesen, and Farah Magrabi. 2021. Realizing AI in Healthcare: Challenges Appearing in the Wild. In Extended Abstracts of the 2021 CHI Conference on Human Factors in Computing Systems (, Yokohama, Japan,) (CHI EA ’21). Association for Computing Machinery, New York, NY, USA, Article 108, 5 pages. https://doi.org/10.1145/3411763.3441347
- Perrier et al. (2015) Trevor Perrier, Nicola Dell, Brian DeRenzi, Richard Anderson, John Kinuthia, Jennifer Unger, and Grace John-Stewart. 2015. Engaging Pregnant Women in Kenya with a Hybrid Computer-Human SMS Communication System. In Proceedings of the 33rd Annual ACM Conference on Human Factors in Computing Systems (Seoul, Republic of Korea) (CHI ’15). Association for Computing Machinery, New York, NY, USA, 1429–1438. https://doi.org/10.1145/2702123.2702124
- Pfeifer Vardoulakis et al. (2012) Laura Pfeifer Vardoulakis, Amy Karlson, Dan Morris, Greg Smith, Justin Gatewood, and Desney Tan. 2012. Using mobile phones to present medical information to hospital patients. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Austin, Texas, USA) (CHI ’12). Association for Computing Machinery, New York, NY, USA, 1411–1420. https://doi.org/10.1145/2207676.2208601
- Prasad et al. (2020) Manya Prasad, Sumit Malhotra, Mani Kalaivani, Praveen Vashist, and Sanjeev K Gupta. 2020. Gender differences in blindness, cataract blindness and cataract surgical coverage in India: a systematic review and meta-analysis. British Journal of Ophthalmology 104, 2 (2020), 220–224.
- Radionova et al. (2023) Natalia Radionova, Eylem Ög, Anna-Jasmin Wetzel, Monika A Rieger, and Christine Preiser. 2023. Impacts of Symptom Checkers for Laypersons’ Self-diagnosis on Physicians in Primary Care: Scoping Review. Journal of Medical Internet Research 25 (2023), e39219.
- Rajashekar et al. (2024) Niroop Channa Rajashekar, Yeo Eun Shin, Yuan Pu, Sunny Chung, Kisung You, Mauro Giuffre, Colleen E Chan, Theo Saarinen, Allen Hsiao, Jasjeet Sekhon, Ambrose H Wong, Leigh V Evans, Rene F. Kizilcec, Loren Laine, Terika Mccall, and Dennis Shung. 2024. Human-Algorithmic Interaction Using a Large Language Model-Augmented Artificial Intelligence Clinical Decision Support System. In Proceedings of the CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿Honolulu¡/city¿, ¡state¿HI¡/state¿, ¡country¿USA¡/country¿, ¡/conf-loc¿) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 442, 20 pages. https://doi.org/10.1145/3613904.3642024
- Roller et al. (2020) Stephen Roller, Emily Dinan, Naman Goyal, Da Ju, Mary Williamson, Yinhan Liu, Jing Xu, Myle Ott, Kurt Shuster, Eric M Smith, et al. 2020. Recipes for building an open-domain chatbot. arXiv preprint arXiv:2004.13637 (2020).
- Seitz et al. (2022) Lennart Seitz, Sigrid Bekmeier-Feuerhahn, and Krutika Gohil. 2022. Can we trust a chatbot like a physician? A qualitative study on understanding the emergence of trust toward diagnostic chatbots. International Journal of Human-Computer Studies 165 (2022), 102848. https://doi.org/10.1016/j.ijhcs.2022.102848
- Singh et al. (2017) Kamaljeet Singh, Arshi Misbah, Pranav Saluja, and Arun Kumar Singh. 2017. Review of manual small-incision cataract surgery. Indian journal of ophthalmology 65, 12 (2017), 1281.
- Skeels et al. (2010) Meredith M. Skeels, Kenton T. Unruh, Christopher Powell, and Wanda Pratt. 2010. Catalyzing Social Support for Breast Cancer Patients. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Atlanta, Georgia, USA) (CHI ’10). Association for Computing Machinery, New York, NY, USA, 173–182. https://doi.org/10.1145/1753326.1753353
- Stamenova et al. (2020) Vess Stamenova, Payal Agarwal, Leah Kelley, Jamie Fujioka, Megan Nguyen, Michelle Phung, Ivy Wong, Nike Onabajo, R Sacha Bhatia, and Onil Bhattacharyya. 2020. Uptake and patient and provider communication modality preferences of virtual visits in primary care: a retrospective cohort study in Canada. BMJ open 10, 7 (2020), e037064.
- Statista (2022) Statista. 2022. Most used social media and messaging apps in India 2022. https://www.statista.com/statistics/1235799/india-most-used-social-media-and-messaging-apps/ Accessed: 2024-02-01.
- Tan et al. (2024) Ting Fang Tan, Kabilan Elangovan, Liyuan Jin, Yao Jie, Li Yong, Joshua Lim, Stanley Poh, Wei Yan Ng, Daniel Lim, Yuhe Ke, et al. 2024. Fine-tuning Large Language Model (LLM) Artificial Intelligence Chatbots in Ophthalmology and LLM-based evaluation using GPT-4. arXiv preprint arXiv:2402.10083 (2024).
- Tang and Lansky (2005) Paul C Tang and David Lansky. 2005. The missing link: bridging the patient-provider health information gap. Health Aff. (Millwood) 24, 5 (Sept. 2005), 1290–1295.
- Tang and Newcomb (1998) P C Tang and C Newcomb. 1998. Informing patients: a guide for providing patient health information. J. Am. Med. Inform. Assoc. 5, 6 (Nov. 1998), 563–570.
- Thieme et al. (2023a) Anja Thieme, Maryann Hanratty, Maria Lyons, Jorge Palacios, Rita Faia Marques, Cecily Morrison, and Gavin Doherty. 2023a. Designing Human-centered AI for Mental Health: Developing Clinically Relevant Applications for Online CBT Treatment. ACM Trans. Comput.-Hum. Interact. 30, 2, Article 27 (mar 2023), 50 pages. https://doi.org/10.1145/3564752
- Thieme et al. (2023b) Anja Thieme, Aditya Nori, Marzyeh Ghassemi, Rishi Bommasani, Tariq Osman Andersen, and Ewa Luger. 2023b. Foundation Models in Healthcare: Opportunities, Risks & Strategies Forward. In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems (, Hamburg, Germany,) (CHI EA ’23). Association for Computing Machinery, New York, NY, USA, Article 512, 4 pages. https://doi.org/10.1145/3544549.3583177
- Tongue et al. (2005) John R. Tongue, Howard R. Epps, and Laura L. Forese. 2005. Communication skills for patient-centered care : Research-based, easily learned techniques for medical interviews that benefit orthopaedic surgeons and their patients. Journal of Bone and Joint Surgery, American Volume 87 (2005), 652–658. https://api.semanticscholar.org/CorpusID:71035586
- Tseng et al. (2023) Yuan-Chi Tseng, Weerachaya Jarupreechachan, and Tuan-He Lee. 2023. Understanding the Benefits and Design of Chatbots to Meet the Healthcare Needs of Migrant Workers. 7, CSCW2, Article 315 (oct 2023), 34 pages. https://doi.org/10.1145/3610106
- Tuli et al. (2018) Anupriya Tuli, Shaan Chopra, Neha Kumar, and Pushpendra Singh. 2018. Learning from and with Menstrupedia: Towards Menstrual Health Education in India. Proc. ACM Hum.-Comput. Interact. 2, CSCW, Article 174 (nov 2018), 20 pages. https://doi.org/10.1145/3274443
- Wang et al. (2020) Ding Wang, Santosh D. Kale, and Jacki O’Neill. 2020. Please Call the Specialism: Using WeChat to Support Patient Care in China. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3313831.3376274
- Wang et al. (2021b) Dakuo Wang, Liuping Wang, Zhan Zhang, Ding Wang, Haiyi Zhu, Yvonne Gao, Xiangmin Fan, and Feng Tian. 2021b. “Brilliant AI Doctor” in Rural Clinics: Challenges in AI-Powered Clinical Decision Support System Deployment. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿Yokohama¡/city¿, ¡country¿Japan¡/country¿, ¡/conf-loc¿) (CHI ’21). Association for Computing Machinery, New York, NY, USA, Article 697, 18 pages. https://doi.org/10.1145/3411764.3445432
- Wang et al. (2021a) Liuping Wang, Dakuo Wang, Feng Tian, Zhenhui Peng, Xiangmin Fan, Zhan Zhang, Mo Yu, Xiaojuan Ma, and Hongan Wang. 2021a. CASS: Towards Building a Social-Support Chatbot for Online Health Community. Proc. ACM Hum.-Comput. Interact. 5, CSCW1, Article 9 (apr 2021), 31 pages. https://doi.org/10.1145/3449083
- Wei et al. (2024) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Brian Ichter, Fei Xia, Ed H. Chi, Quoc V. Le, and Denny Zhou. 2024. Chain-of-thought prompting elicits reasoning in large language models. In Proceedings of the 36th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’22). Curran Associates Inc., Red Hook, NY, USA, Article 1800, 14 pages.
- Weisz et al. (2019) Justin D. Weisz, Mohit Jain, Narendra Nath Joshi, James Johnson, and Ingrid Lange. 2019. BigBlueBot: teaching strategies for successful human-agent interactions. In Proceedings of the 24th International Conference on Intelligent User Interfaces (Marina del Ray, California) (IUI ’19). Association for Computing Machinery, New York, NY, USA, 448–459. https://doi.org/10.1145/3301275.3302290
- Wilcox et al. (2010) Lauren Wilcox, Dan Morris, Desney Tan, and Justin Gatewood. 2010. Designing Patient-Centric Information Displays for Hospitals. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Atlanta, Georgia, USA) (CHI ’10). Association for Computing Machinery, New York, NY, USA, 2123–2132. https://doi.org/10.1145/1753326.1753650
- Williams and Moser (2019) Michael Williams and Tami Moser. 2019. The art of coding and thematic exploration in qualitative research. International Management Review 15, 1 (2019), 45–55.
- Wutz et al. (2023) Maximilian Wutz, Marius Hermes, Vera Winter, and Juliane Köberlein-Neu. 2023. Factors influencing the acceptability, acceptance, and adoption of conversational agents in health care: integrative review. Journal of Medical Internet Research 25 (2023), e46548.
- Xiao et al. (2023) Ziang Xiao, Q. Vera Liao, Michelle Zhou, Tyrone Grandison, and Yunyao Li. 2023. Powering an AI Chatbot with Expert Sourcing to Support Credible Health Information Access. In Proceedings of the 28th International Conference on Intelligent User Interfaces (Sydney, NSW, Australia) (IUI ’23). Association for Computing Machinery, New York, NY, USA, 2–18. https://doi.org/10.1145/3581641.3584031
- Yadav et al. (2019) Deepika Yadav, Prerna Malik, Kirti Dabas, and Pushpendra Singh. 2019. Feedpal: Understanding Opportunities for Chatbots in Breastfeeding Education of Women in India. Proc. ACM Hum.-Comput. Interact. 3, CSCW, Article 170 (nov 2019), 30 pages. https://doi.org/10.1145/3359272
- Yang et al. (2024) Ziqi Yang, Xuhai Xu, Bingsheng Yao, Ethan Rogers, Shao Zhang, Stephen Intille, Nawar Shara, Guodong Gordon Gao, and Dakuo Wang. 2024. Talk2Care: An LLM-based Voice Assistant for Communication between Healthcare Providers and Older Adults. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 8, 2, Article 73 (may 2024), 35 pages. https://doi.org/10.1145/3659625
- Ye et al. (2020) Qunru Ye, Yanxian Chen, William Yan, Wei Wang, Jingxian Zhong, Cong Tang, Andreas Müller, Bo Qiu, et al. 2020. Female gender remains a significant barrier to access cataract surgery in South Asia: A systematic review and meta-analysis. Journal of ophthalmology 2020 (2020).
- Yeh et al. (2022) Su-Fang Yeh, Meng-Hsin Wu, Tze-Yu Chen, Yen-Chun Lin, XiJing Chang, You-Hsuan Chiang, and Yung-Ju Chang. 2022. How to Guide Task-oriented Chatbot Users, and When: A Mixed-methods Study of Combinations of Chatbot Guidance Types and Timings. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (New Orleans, LA, USA) (CHI ’22). Association for Computing Machinery, New York, NY, USA, Article 488, 16 pages. https://doi.org/10.1145/3491102.3501941
- Yilmaz and Dogan (2024) Ibrahim Edhem Yilmaz and Levent Dogan. 2024. Talking technology: exploring chatbots as a tool for cataract patient education. Clinical and Experimental Optometry (2024), 1–9.
- YM et al. (2023) Pradyumna YM, Vinod Ganesan, Dinesh Kumar Arumugam, Meghna Gupta, Nischith Shadagopan, Tanay Dixit, Sameer Segal, Pratyush Kumar, Mohit Jain, and Sriram Rajamani. 2023. PwR: Exploring the Role of Representations in Conversational Programming. arXiv:2309.09495 [cs.HC]
- Zhang (2010) Yan Zhang. 2010. Contextualizing consumer health information searching: an analysis of questions in a social Q&A community. In Proceedings of the 1st ACM international health informatics symposium. 210–219.
Appendix A Appendix
A.1. Cataract Bot Implementation Details
A.1.1. Cataract Bot LLM Prompts
The CataractBot system leverages LLM (GPT-4 in our case) for these four tasks (Table 3):
Response Generation: For every medical and logistical question asked by the patient/attendant, the system performs a vector search on the Knowledge Base (KB) to extract the three most relevant data chunks related to the query. The LLM is then prompted (Table 3) to extract an answer for the query from these data chunks. Note: Question Classification is part of the same LLM call to improve efficiency.
Related Questions Generation: For every medical and logistical question, the system prompts LLM to generate three related questions based on the preceding query. Note: The 72 character limit is due to the WhatsApp’s message limit in interactive suggestions.
Final Response Generation: If the expert marks the initial LLM-generated response as incorrect or incomplete, the system prompts LLM to generate the final response by merging both the initial response and the expert’s edit suggestion.
Shorten Response: If the generated response exceeds WhatsApp’s message limit of 700 characters, the system prompts LLM to summarize it within the specified character limit.
| System Prompt | Query Prompt | |||||||||||||||||||||||||||||||||||||||||||
A.2. Chatbot Usability Form
Responses were given on a 5-point Likert scale, ranging from Strongly Disagree to Strongly Agree.
CataractBot understands me well.
CataractBot’s responses are easy to understand.
CataractBot’s responses were useful, appropriate, and informative.
CataractBot responds quickly.
CataractBot seems to have a good grasp of medical knowledge.
CataractBot is kind and helpful.
CataractBot is easy to use.
I would be willing to use CataractBot (or a similar bot before/after a major surgical treatment) in future.