### v5

# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

Abstract.
The healthcare landscape is evolving, with patients seeking reliable information about their health conditions and available treatment options. Despite the abundance of information sources, the digital age overwhelms individuals with excess, often inaccurate information. Patients primarily trust medical professionals, highlighting the need for expert-endorsed health information. However, increased patient loads on experts has led to reduced communication time, impacting information sharing. To address this gap, we developed CataractBot [^1] Our code is publicly available at https://github.com/microsoft/BYOeB.. CataractBot answers cataract surgery related questions instantly using an LLM to query a curated knowledge base, and provides expert-verified responses asynchronously. It has multimodal and multilingual capabilities. In an in-the-wild deployment study with 49 patients and attendants, 4 doctors, and 2 patient coordinators, CataractBot demonstrated potential, providing anytime accessibility, saving time, accommodating diverse literacy levels, alleviating power differences, and adding a privacy layer between patients and doctors. Users reported that their trust in the system was established through expert verification. Broadly, our results could inform future work on designing expert-mediated LLM bots.
GPT-4, Generative AI, LLM, Question Answering Bot, Medical, Healthcare, Surgery
$c$s: Human-centered computing Interaction designccs: Applied computing Health care information systemsccs: Human-centered computing Ubiquitous and mobile computing systems and tools
1. Introduction
The evolving landscape of healthcare witnesses a significant shift with patients assuming more proactive roles in their care journeys (Klasnja et al., 2010a; Kinnersley et al., 2007), resulting in an increased demand for information. Patients and their caregivers seek accessible, comprehensive, and reliable information about their symptoms, diagnoses, treatment options, potential risks, and preventive measures (Christalle et al., 2019). To satisfy these information needs, patients actively explore diverse sources, including online resources, friends and family, support groups, and direct communication with healthcare professionals (Skeels et al., 2010). In particular, for major treatments like surgery, satisfying the demand for information becomes crucial (Christalle et al., 2019), with questions spanning pre-, during-, and post-treatment phases. Studies highlight the anxiety patients and caregivers experience regarding such treatments (Klasnja et al., 2010b), emphasizing the correlation between anxiety and negative clinical outcomes (Byrne and Heyman, 1997; Dahlén and Janson, 2002). Similarly, access to information has been found to significantly reduce anxiety (Wilcox et al., 2010) and improve patient satisfaction and clinical outcomes (Kurtz, 2002; Wilcox et al., 2010), thus underlining its pivotal role.
Despite the abundance of available information sources, the digital age often overwhelms individuals with an excess of information, much of which is inaccurate or unreliable (Clarke et al., 2016a; Kinnersley et al., 2007). This poses a challenge in discerning trustworthy sources from misleading ones. Patients also frequently encounter difficulties finding information online at an appropriate level, ranging from oversimplified to overly technical (Xiao et al., 2023). As a result, patients tend to rely on the experiences of friends and family members who have undergone similar treatments, although identifying them can be challenging (Civan-Hartzler et al., 2010). Therefore, patients primarily trust their doctor and the hospital staff responsible for their treatment, turning to them for any medical or logistical queries (Attfield et al., 2006; Clarke et al., 2016a). Studies examining patients’ information needs reveal a strong desire for doctor-endorsed health information (Tang and Newcomb, 1998).
However, the escalating pressure on doctors to accommodate more patients has led to reduced time per patient, affecting communication and information sharing. In developing nations, a lower doctor-to-patient ratio hinders personalized attention and comprehensive guidance (Wang et al., 2021b), while in developed countries, managing Electronic Health Records (EHR) often diminishes direct patient communication quality and available time (Cajander and Grünloh, 2019a). Previous studies (Tongue et al., 2005; Kinnersley et al., 2007) have found that doctors often underestimate patients’ information needs and overestimate the amount of information they provide. However, providing all potentially relevant information risks overwhelming patients with cognitive overload (Christalle et al., 2019). Surveys consistently highlight patients’ desire for improved communication with their healthcare providers (Duffy et al., 2004). This information exchange extends beyond doctor-patient interactions to encompass relationships between various medical professionals and patients, such as patient-nurse interactions (Lotfi et al., 2019; Wang et al., 2020). Moreover, enhanced access to information not only benefits patients but also has the potential to reduce unnecessary visits, alleviating the burden on doctors and medical staff. To summarize, as Tang and Lansky (2005) correctly stated: “Patients have little access to information and knowledge that can help them participate in, let alone guide, their own care… A simple, non-urgent exchange of questions and answers is often all that is required.”.
To address patients’ information needs, particularly during surgical treatments, we propose a chatbot solution integrated with medical professionals to provide accessible, reliable and comprehensive responses. Leveraging the capabilities of large-language models (LLMs) and the widespread use of smartphones and instant messaging services such as WhatsApp, this chatbot serves as a 24/7 resource capable of understanding intricate human queries and providing accurate information. In collaboration with a tertiary eye hospital in Bangalore, India, we exemplified this experts-in-the-loop chatbot approach by developing CataractBot, which answers queries related to cataract surgery from patients and their attendants. Recognizing the limitations of generic LLMs in the medical domain, we utilized Retrieval-Augmented Generation (Lewis et al., 2020) with an LLM over a custom knowledge base curated by doctors and hospital staff to provide hospital-specific and culturally sensitive responses. CataractBot has several features to enhance adoption and sustained use. Capitalizing on the widespread adoption of smartphones, we chose a mobile-first design paradigm, deploying the chatbot entirely on the ubiquitous WhatsApp platform. To cater to diverse hospital visitors with varying literacy levels and technological proficiency, we made CataractBot multilingual (supporting five languages, including English, Hindi, and Kannada) and multimodal (accepting both speech and text inputs). We use an LLM to process complex and ill-formed queries, generating instant responses from the curated knowledge base. Crucially, CataractBot incorporates a system where the LLM’s responses are asynchronously verified by experts and corrected if needed. Ophthalmologists verify answers to medical queries and patient coordinators to logistical ones. These expert-provided edits are used to update the knowledge base, to minimize future expert intervention.
Our work aims to answer the following research questions: (RQ1) What is the role of CataractBot in meeting the information needs of patients (and attendants) undergoing surgery? (RQ2) How do the different features of CataractBot, including LLM-generated answers, experts-in-the-loop and multimodality, contribute to its usage? (RQ3) How can CataractBot be integrated into the current doctor-patient workflow? To answer these questions, we conducted an in-the-wild deployment study involving multiple stakeholders (See Table 1): 49 information seekers (19 patients, 30 attendants) and 6 experts (4 doctors and 2 patient coordinators). Our findings indicated that patients and attendants appreciated CataractBot because it alleviated hesitations associated with power differences in asking questions, and accommodated individuals with low literacy and tech proficiency through multilingual and multimodal support. Patients and attendants reported that their trust in the system was established through the verification performed by doctors and coordinators. Experts praised the bot as a facilitator, introducing a privacy layer between them and patients, and providing flexibility through asynchronous communication.
Our main contributions are: (1) The design and development of a novel LLM-powered, experts-in-the-loop, multilingual, multimodal chatbot, created in collaboration with doctors and hospital staff to assist cataract surgery patients with their information needs. (2) The deployment of our bot in a real-world multi-stakeholder setting, among 49 cataract surgery patients/attendants, four doctors, and two patient coordinators at a tertiary eye hospital in India. (3) Drawing from a comprehensive mixed-methods analysis of gathered data, the paper offers pivotal insights specific to CataractBot and broader implications of LLM-powered experts-in-the-loop chatbots.
2. Related Work
Our work is primarily informed by two areas of relevant research: solutions addressing patients’ information needs and the use of conversational agents, particularly those powered by LLMs, in healthcare settings.
2.1. Patients’ Information Needs
Patients and their attendants seek information throughout the various stages of their medical journey, such as preparing for doctor consultations (Radionova et al., 2023; Leong et al., 2018), verifying diagnoses and treatment plans (Eriksson-Liebon et al., 2021; Bickmore et al., 2009; Christalle et al., 2019), assessing the need for clinical interventions (Yang et al., 2024; Tseng et al., 2023), and aiding in the recovery process (Wang et al., 2021a; Civan-Hartzler et al., 2010; Fadhil, 2018). Studies emphasize that satisfying these diverse information needs improves patient understanding, motivates adherence to treatment regimens, and contributes to overall satisfaction with healthcare services (Tang and Newcomb, 1998). Patients primarily trust doctors with their medical queries (Ding et al., 2020). However, traditional in-person consultations pose difficulties, including high costs (Bodenheimer and Smith, 2013), long wait times, and a strong hierarchy in the patient-physician relationship (Fochsen et al., 2006). Patients often miss opportunities to ask additional (clarification) questions during these brief encounters, as they face information overload (Ferré et al., 2020; Kinnersley et al., 2007). Doctors also struggle with time due to staff shortages straining healthcare ecosystems (Bodenheimer and Smith, 2013; Akerkar and Bichile, 2004). Thus, patients and caregivers are increasingly turning to the internet for easier access to medical information (Xiao et al., 2023; Chandwani and Kulkarni, 2016). However, the challenge lies in identifying relevant and appropriate information. Individuals struggle to understand complex health information or may even consume inaccurate information, making the process frustrating (Zhang, 2010; Arora et al., 2008; Clarke et al., 2016b).
Information seekers may also consult experienced patients within their social networks who can share insights related to managing similar health conditions (Civan-Hartzler et al., 2010; Skeels et al., 2010). In the absence of such known individuals, people may resort to social media, which is not only ineffective for quality health support due to widespread misinformation and poor management (Malki et al., 2024), but also risky in terms of personal privacy (De Choudhury et al., 2014). Chandwani and Kulkarni (2016) found that although most doctors regard the practice of seeking health information online as an affront to their authority, some see it as inevitable–and even as an opportunity to elevate patients’ health literacy through technology.
Recent work in the field of Computer-Human Interaction has explored novel avenues to address these medical information gaps. For instance, Leong et al. (2018) propose interactive ‘science museum’ exhibits in clinical waiting rooms to educate children and their parents about sickle cell anemia, Wilcox et al. (2010) and Bickmore et al. (2009) propose information displays in hospital rooms to provide patients with real-time status updates and treatment details, and Pfeifer Vardoulakis et al. (2012) propose using mobile phones to present dynamic, interactive reports on patients’ progress and care plans throughout their emergency department stay. Although these are specific solutions for certain diseases and infrastructures, discussions have been ongoing around making Electronic Health Record (EHR) data accessible to patients (Cajander and Grünloh, 2019b; Tang and Lansky, 2005; Chung, 2017). Patient-facing EHR services have received mixed reviews, with healthcare professionals expressing concerns about adding strain to already understaffed healthcare systems, while patients generally welcome the idea (Cajander and Grünloh, 2019b).
Prior work has also explored “chat” as an interface to address patients’ information needs, by connecting them with doctors in both synchronous (Li et al., 2023) and asynchronous (Johansson et al., 2020a, b) manners. Synchronous messaging requires both patient and doctor to be simultaneously active, allowing instant feedback and coherent dialogue. On the other hand, asynchronous messaging platforms like WhatsApp (Yadav et al., 2019), WeChat (Wang et al., 2020), SMS (Perrier et al., 2015), email (Makarem and Antoun, 2016), or web portals (Mamykina et al., 2008; Tuli et al., 2018) enable communication at the convenience of both parties. Research indicates that both doctors and patients prefer asynchronous texts over instant audio/video calls (Stamenova et al., 2020; Johansson et al., 2020a), despite lacking instant feedback (Li et al., 2023).
In summary, patients need reliable, relevant, and timely information. However, obtaining this from busy doctors and hospital staff is challenging. This dichotomy results in patients either receiving generic, unverified, and hard-to-understand information from the internet and social circles, or, if lucky, obtaining personalized, verified answers from healthcare experts—at a higher cost. Our solution aims to strike a balance by providing a generic, high-quality LLM answer, later verified by a medical professional with minimal increase to their workload.
2.2. LLM Chatbots in Healthcare
Chatbots enable natural language conversations, offering a minimal learning curve and a personalized experience (Thirunavukarasu et al., 2023). AI-powered chatbots surpass rule-based platforms in language understanding, allowing users to build complex queries message by message while retaining context (Xiao et al., 2023). With the advent of LLMs, building complex chatbots has become significantly easier, enabling researchers to apply LLM-powered bots across diverse healthcare settings (Thirunavukarasu et al., 2023), including pre-consultation data collection (Li et al., 2024), chronic disease management (Montagna et al., 2023; Hao et al., 2024), and mental health support (Kumar et al., 2023; Jo et al., 2023; Kim et al., 2024). Commercial solutions, such as Babylon Health (eMed, 2024), now provide individuals with instant access to any health-related information. However, LLM-powered healthcare tools have generally been criticized for hallucinations, errors, and a lack of transparency in their reasoning (Au Yeung et al., 2023; Denecke et al., 2024). As the involvement of doctors in these bots is–at most–limited to the data curation step (Xiao et al., 2023), they are susceptible to generating inaccurate information and are “not ready for clinical use” (Au Yeung et al., 2023).
Several methods have been proposed to address these issues, including chain-of-thought prompting (Wei et al., 2024), retrieval-augmented generation (Lewis et al., 2020), and few-shot learning. Seitz et al. (2022) identified that patients’ trust in healthcare chatbots is primarily based on a rational evaluation of their capabilities, while trust in human medical professionals relies more on emotional connections and personal rapport. Integrating doctors into the loop with automated LLM chatbots could combine the best of both worlds, a concept we propose in CataractBot. In our work, we leverage LLM-based responses in real-time, while human medical experts verify and respond asynchronously. This hybrid approach is a novel and reliable method for addressing healthcare information needs.
We chose cataract surgery–one of the world’s most common procedures (McGhee et al., 2020)–as the use case for developing and testing our system. This decision was motivated by our longstanding collaboration with Sankara Eye Hospital, an eye hospital in India. Also, cataract surgery involves well-defined protocols and common patient queries, making it an ideal starting point for testing the bot’s ability to deliver accurate, standardized information from a curated knowledge base. However, our solution is versatile, and we believe it can effectively support a range of health conditions, including mental health and chronic care. Given our specific focus on cataract surgery, we examined literature at the intersection of ophthalmology and LLMs. Bernstein et al. (2023) reported that responses generated by LLMs to patient eye-care-related questions are comparable to those written by ophthalmologists. Based on similar findings, Ittarat et al. (2023) proposed integrating chatbots into ophthalmology practices to provide 24/7 support for common inquiries on eye conditions. Additionally, Yilmaz and Dogan (2024) reported that chatbots are an effective tool for educating cataract surgery patients. Our work contributes to this body of literature.
Finally, real-world healthcare deployments of AI are limited due to socio-technical uncertainties in the last mile (Osman Andersen et al., 2021). To avoid such shortcomings, as recommended by Thieme et al. (2023b, a), we designed and studied our system within a tertiary eye hospital in India and examined its integration into clinical workflows.
3. Formative Study
[…]
3.1. Background
Our design requirements were shaped by the workflows and information needs specific to cataract surgery at Sankara Eye Hospital. In the current workflow, patients (with their attendants) first consult a doctor who diagnoses the cataract’s maturity and type, recommending the urgency of surgery and potential complications. After deciding on surgery, the patient and attendant meet with a patient coordinator for guidance on pre- and post-operative measures, lens options across different budget ranges, and logistical tasks such as insurance and scheduling. Patients also undergo pre-operative medical tests. For further questions, they can either visit the hospital to meet the doctor or contact the coordinator by phone. On the day of surgery, patients arrive with their attendants in the morning, undergo surgery by mid-morning, and are discharged by the afternoon or evening. The doctor and coordinator provide post-surgery care instructions, covering medication, exercise, bathing, travel, and screen time. One week post-surgery, patients return for a follow-up appointment with the operating doctor.
3.2. Design Requirements
Accuracy. The chatbot’s ability to answer both medical and logistical questions with accuracy, precision and contextual relevance was found to be essential given the medical context. While a standard LLM like GPT-4 (OpenAI et al., 2023) might offer accurate information, it may not be tailored to India- or hospital-specific nuances. For instance, doctors informed us that although the phacoemulsification method is prevalent globally for cataract surgeries, they often favor manual small incision-based procedures (Singh et al., 2017) due to their cost-effectiveness. Similarly, cultural considerations significantly influence pre- and post-surgery questions. As a doctor stated, “People have unique questions… like ‘when can I do nasal yoga post-op?’.” Further, queries related to diet require awareness of Indian food options for precise guidance. Thus, our bot relies on a custom knowledge base curated by hospital staff.
Trustworthiness. Access to trustworthy information is crucial during medical treatment. As one doctor noted, “The success of treatment is not only in medicine, but the trust they place on us.” Patients and their attendants typically place their utmost trust in the operating doctor and hospital staff. Even when patients learn something from the internet or other sources, they often confirm that information with their doctor. To ensure trustworthiness, we decided that (a) each response would be verified and, if necessary, edited by a doctor or patient coordinator, and (b) patients (or their attendants) would be explicitly notified upon expert verification. Note: We intentionally designed CataractBot to be non-specific to individual patients by not connecting it with the hospital’s scheduling system or EHR data. This decision was made because LLMs can memorize training data, including personally identifiable information (PII) like emails and phone numbers, and leak it during inference (Borkar, 2023), raising privacy concerns and potentially breaking users’ trust.
Timeliness. Discussions with hospital staff highlighted the crucial need for real-time responses. Patient coordinators noted that patients frequently reach out to them over multiple phone calls to ask logistical questions and relay medical queries to their consulting doctors. Due to experts’ busy schedules, immediate responses are not always feasible. This often results in patient seeking information from other sources (including the internet), which can propagate misinformation, or fail to completely address their information needs. To counter this, we decided that the chatbot should provide an instantaneous response to every patient query using the LLM and a custom knowledge base. Subsequently, these responses get reviewed and refined by doctors/coordinators. Moreover, doctors mentioned their unavailability for several hours while performing surgeries. To ensure swift verification, CataractBot forwards queries to an ‘escalation’ expert if they remains unverified for 3 hours.
Usability. After cataract surgery, patients can reference their discharge summaries for any queries, but this is challenging due to their temporarily reduced vision. Prior research (Grönvall and Verdezoto, 2013) has identified the crucial role family members play in monitoring the patient’s health. Coupled with doctors’ input (“Patients ask hardly 10% of the question… attendants are the ones who ask.”), this underscored the necessity for the bot to be usable by both the patients and their attendants. To cater to diverse backgrounds, we made the chatbot multilingual and multimodal. We chose WhatsApp as the messaging platform due to its widespread use in India (Statista, 2022) and therefore minimal learning curve and quick onboarding without additional installations, despite the richer feature set of other platforms like Telegram or the development flexibility of a custom chatbot. Considering doctors’ demanding schedules, often involving consulting $\sim$50 patients and performing $\sim$10 cataract surgeries a day, we implemented features to minimize their bot-related workload: (1) one-click interaction to verify an answer, (2) allow experts to provide corrections using informal messages (Beasley, 2009), (3) handle spelling and grammar errors, and (4) use expert edits to enhance the bot’s knowledge base, minimizing similar edits in future.
4. CataractBot System Design
Here, we describe the key components of the CataractBot system (Figure 2)–input language and modality, response generation and verification, escalation and reminders–and provide a detailed account of its implementation.
Figure 1. CataractBot provides an initial response to the patient/attendant by querying the knowledge base. The doctor (or coordinator, if the question is logistical) verifies and corrects this response, and the patient/attendant is notified.:
Table 1. Stakeholders and their roles in the CataractBot socio-technical system
Stakeholder Role
Information seeker Patient
Person scheduled for cataract surgery.
Asks CataractBot surgery-related questions.
Attendant
Person accompanying the patient (e.g., child of the patient).
Asks CataractBot surgery-related questions.
Expert Doctor
Operating
doctor
Surgeon scheduled to operate on the patient.
Verifies CataractBot’s answers to users’ medical questions.
Escalation
doctor
Senior surgeon. Verifies CataractBot’s answers to medical
questions that the operating doctor is unable to address in time.
Knowledge
base expert
Senior surgeon. Selects and edits verified answers for
addition to CataractBot’s knowledge base.
Patient coordinator
Operating
coordinator
Liaison between patients/attendants and operating doctors.
Verifies CataractBot’s answers to users’ logistical questions.
Escalation
coordinator
Senior coordinator. Verifies CataractBot’s answers to logistical
questions that the operating coordinator is unable to address in time.
4.1. Components
4.1.1. Input Language and Modality
Sankara Eye Hospital in Bangalore (in the state of Karnataka), referred to as the Silicon Valley of India, caters to patients from various linguistic, educational, and technical backgrounds, including those from the Information Technology sector and neighboring states. Analysis of a 2011 census highlights Bangalore as one of India’s most linguistically diverse cities (Office of the Registrar General & Census Commissioner, 2011). To accommodate this varying patient demography, CataractBot is designed to support five languages: English, Hindi (the local language of nine states in India), Kannada (the local language of Karnataka), and Tamil and Telugu (the local languages of two neighboring states of Karnataka). As the expert may not be proficient in all five languages, their interactions with the bot are exclusively in English. Upon onboarding, the language preference of the patient and their attendant is collected through an online form, and CataractBot initiates the conversation by sending a set of ‘welcome messages’ in the chosen language. Among these messages is an option allowing users to modify their preferred language in future. To engage with users with different levels of literacyCataractBot supports both text and speech inputs. For every voice message, CataractBot sends both a text response and a spoken version of the text as an audio message.
Recognizing the challenge users face in initiating a conversation, (‘Blank Page Syndrome’ (Moden, 2022; YM et al., 2023)), CataractBot provides a set of three frequently asked questions along with the welcome message. Additionally, every response includes a “What to do next?” prompt, offering three related questions based on the preceding query (Figure 4F in Appendix A.1.2). Users can select one of the suggested questions or pose a different query to continue the conversation. This feature lowers the barrier to entry (as formulating questions has been found to be non-trivial for low-literate individuals (Tseng et al., 2023)), enhances the natural flow of conversation, and saves time.
4.1.2. Response Generation, Verification and Icons
Upon receiving a message, CataractBot classifies it as a medical question (e.g., dos and donts before and after surgery), logistical question (e.g., scheduling or insurance related), small talk (e.g., greetings or expressions of gratitude), or ‘other’, and provides a response in real-time. For medical and logistical questions, the bot strictly employs the knowledge base curated by the hospital’s medical team to generate an appropriate response. This custom knowledge base comprises of various cataract surgery and hospital-specific documents curated by hospital staff (doctors, patient coordinators, and members of Quality Control and the Patient Safety team). These documents, comprising approximately 30 pages in total, include the Consent Form, Standard Operating Procedures, FAQs, Pre- and Post-Operative Guidelines, and Hospital Information. Some documents, like Post-Operative Guidelines, were included in patients’ discharge summaries. Others, such as SOPs and FAQs, existed in the hospital ecosystem but were not directly accessible to patients. The generated response includes a question mark icon [Uncaptioned image] as a ‘reaction’ indicating the unverified status (Figure 4E in Appendix A.1.2). In instances where the custom knowledge base lacks an answer, the bot responds with a template “I don’t know” response. For small talk messages (such as “Hello” or “Thank you for the information”), the chatbot provides corresponding small talk responses.
For medical questions, the patient’s operating doctor (Table 1) receives a message comprising of the question asked, the bot’s response, and patient demographics (Figure 1A). As Rajashekar et al. (2024) found that citations improve trust in LLM-generated responses, this message also includes citations of the documents used to generate the response. The doctor is prompted with the question, “Is the answer accurate and complete?” offering three response options: “Yes”, “No”, and “Send to Patient Coordinator”. Tapping ‘Yes’ replaces the question mark icon [Uncaptioned image] in the information seeker’s received answer with a green tick icon [Uncaptioned image], confirming verification (Figure 1F). Additionally, the bot notifies the information seeker that the answer has been verified, tagging that particular response (Figure 1G). Tapping ‘No’ replaces the question mark [Uncaptioned image] with a red cross icon [Uncaptioned image], indicating an incorrect answer (Figure 1C), and asks the information seeker to await a corrected response from the doctor (Figure 1D). The doctor is asked to provide a correction; they are not required to edit the bot’s message, instead, they can offer a correction in free form text. The system combines the bot’s initial answer with the doctor-provided correction to frame a new response, which is delivered to the user with a green tick icon [Uncaptioned image] (Figure 1E). This new response is not sent to the doctor, to minimize their workload by limiting verification to a single round. If a question is misclassified, i.e., a logistical question is sent to the doctor, they have the option to ‘Send to Patient Coordinator’. Similar to the doctors’ workflow, patient coordinators verify and provide corrections to the bot’s responses for logistical questions. We deliberated between displaying unverified responses in real-time or only verified responses post-expert verification and correction. Through our formative study and discussions with hospital staff, it became clear that providing real-time responses was crucial. Delayed responses might lead to patients making repeated calls to the hospital or resorting to online sources, failing to meet their information needs.
4.1.3. Escalation Mechanism
The CataractBot system employs an intricate escalation and reminder mechanism to ensure swift verification. In the doctor’s WhatsApp interface, unanswered questions are marked with a question mark icon [Uncaptioned image] (Figure 4E in Appendix A.1.2) and answered questions with a green tick [Uncaptioned image], enabling them to easily identify pending queries. If an answer is not verified by the operating doctor within three hours, it is automatically sent to the designated escalation doctor (Table 1). Both the operating and escalation doctors can then verify/correct the response. Selecting ‘Yes’ or ‘Send to Patient Coordinator’ immediately marks the query with a green tick (Figure 1F) for both the doctors, indicating task completion. However, if either doctor selects ‘No’, the green tick appears only after that doctor provides a correction. If neither the operating nor escalation doctor verifies a question within six hours, both receive a reminder notification indicating the pending status. Additionally, every four hours during working hours (at 8 am, 12 pm, and 4 pm), a list of all questions pending verification for more than six hours is sent to both the operating and escalation doctors. This reminder was added based on the doctors’ feedback. This workflow is mirrored for the operating and escalation patient coordinators.
4.1.4. Knowledge Base Update Process
To minimize experts’ labor, we use expert-provided edits to update the knowledge base, which increases the likelihood of ‘Yes’ responses from experts to similar questions in the future. However, certain responses, such as those specific to individual patients (e.g., “When to reach the hospital for my surgery?”) must not be added to the knowledge base as they are not generalizable. We appointed a senior cataract surgeon as the ‘knowledge base expert’ (Table 1), who received a spreadsheet via email at 8pm daily. Their task involves reviewing each row, which consists of question-answer pairs, and determining whether the information should be added to the knowledge base (by responding with a ‘Yes’/‘No’ in the ‘Should Update Knowledge Base?’ column), and if so, modifying as needed the ‘Final Answer for Knowledge Base’ column, containing the bot’s updated answer based on the expert’s correction. At 3 am daily, the system extracts ‘Question’ and ‘Final Answer for Knowledge Base’ data from all rows marked ‘Yes’, and append these to the knowledge base in the ‘expert-FAQ’ document. Additionally, we prompt the LLM to prioritize this ‘expert-FAQ’ document within the knowledge base. This ensures that CataractBot improves with these updates, gradually enhancing its accuracy over time.
4.2. Implementation Details
The CataractBot system relies on these five components (Figure 2): (a) LLM: for response generation, (b) Vector Database: for storing and retrieving the custom knowledge base, (c) Language Technologies: for translation and transcription, (d) WhatsApp Services: for message exchange, and (e) Cloud Storage: for storing interaction logs.
We opted for GPT-4, the leading LLM at the time of our system development in May 2023 (OpenAI et al., 2023). The documents for the custom knowledge base are ingested as data chunks into a Chroma vector database, using the OpenAI model ‘text-embedding-ada-002’ to generate embeddings. Upon receiving a question, GPT-4 classifies the query type. For a medical/logistical question, CataractBot employs a retrieval-augmented generation (Lewis et al., 2020) approach. This involves performing a vector search on the knowledge base to extract the three most relevant data chunks. GPT-4 is then prompted (full prompt available in Appendix A.1.1) to generate an answer for the query from these data chunks. If the answer is not present, the bot responds with an “I don’t know” message.
The prompt underwent several iterations. To validate its effectiveness, a doctor and a patient coordinator recorded 153 questions posed by patients/attendants scheduled for cataract surgery over a week. We used our GPT-4 prompt to obtain answers to these questions from the custom knowledge base, and subsequently the generated responses were evaluated by the same doctor and coordinator. This not only helped refine the LLM prompt, but also led the medical staff to improve existing documents and contribute additional documents to enrich the knowledge base. This process was iterated thrice.
Additionally, we utilize GPT-4 to incorporate corrections provided by doctors/coordinators to generate the final expert corrected response. This involved prompting GPT-4 with the patient’s question, the initial response from the bot, and the expert-provided correction, and asking it to generate a revised answer considering the expert’s input. Our experience showed that GPT-4 executed this task well. Regarding language integration, GPT-4 primarily comprehends and responds in English (Ahuja et al., 2023). To bridge the language gap, our system adhered to the standard approach (ait, [n. d.]) of translating Indic languages (Hindi, Kannada, Tamil, and Telugu) into English for input using Azure AI Translator. It then translated GPT-4’s English responses back into the respective Indic languages for output. To facilitate speech as input and audio as output, CataractBot uses Azure speech-to-text and text-to-speech models. (Note: GPT-4 lacks support for speech input.) Finally, all text and audio interactions between the CataractBot system and users (including patients, attendants, operating and escalation doctors, coordinators, and knowledge base expert) get logged in a cloud storage for further analysis.
Figure 2. Overview of CataractBot System:
5. Study Design
To evaluate the effectiveness of this LLM-powered experts-in-the-loop chatbot in addressing the informational needs of patients undergoing cataract surgery, we conducted a mixed-method user study during Nov 2023-Jan 2024 at our collaborators’ institute, Sankara Eye Hospital. This is one of the leading tertiary eye care and teaching institutions in Bangalore, India. It typically attends to more than 500 patients every day, and conducts over 50 cataract surgeries daily. This study was approved by both the Scientific and Ethics Committees of Sankara Eye Hospital. None of the study participants—patients, attendants, doctors, or coordinators—received any financial incentives for their participation, in accordance with hospital norms.
5.1. Procedure for Patients and Attendants
The operating patient coordinator was tasked to assess the patient’s eligibility for the study based on specific criteria: (a) age of 18 or older, (b) fluent in one of the five languages supported by CataractBot, (c) scheduled for cataract surgery within a week, (d) no history of cataract surgery in the last 6 months (as recent patients would likely have minimal informational needs), and (e) having one of the three participating operating doctors as their surgeon. If these criteria were met, the patient and their attendant were directed by the coordinator to meet a researcher (the first author) stationed at the hospital.
The researcher introduced them to CataractBot and outlined the study’s protocol which involved using the bot for around two weeks and engaging in an interview pre- and post-surgery. If they agreed to participate, participants were requested to sign a consent form, after which the researcher filled a web-based onboarding form. This form included details such as the patient and attendant’s phone numbers linked to WhatsApp, preferred languages, consulting doctor and coordinator, surgery date, and basic demographic information (age, gender, and education). Upon form submission, participants received ‘welcome messages’ from CataractBot. The researcher instructed them to ask a trial question, either by choosing from suggested questions or by typing/speaking in their preferred language. Upon receiving a response, participants were briefed on the icons and expert verification system. Throughout this process, the researcher encouraged them to ask any questions regarding the bot’s usage. After onboarding, participants received two daily reminder messages (at 7:30 am and 4 pm) prompting them to “ask any cataract surgery related questions” until one week post surgery, when their access was revoked.
We conducted two semi-structured interviews: one on the surgery day and another a week post-surgery. These specific days were chosen because they coincided with the patient’s required hospital visits. The interviews explored their overall experience, specific features they liked or disliked, suggestions for improvement, and questions regarding the trustworthiness, timeliness, and accuracy of responses. Both interviews followed a similar structure, differing only in how the bot supported the participant before versus after surgery. After each interview, participants were requested to fill out a chatbot usability form. Those who agreed, rated their chatbot experience through eight questions, using a five-point Likert scale, covering aspects like ease of use, understandability, and timeliness (Appendix A.3). These questions were adapted from metrics outlined in prior usability evaluations of healthcare bots (Abd-Alrazaq et al., 2020; Holmes et al., 2023). Patient-attendant pairs were interviewed together but filled separate usability forms.
All interviews with patients and attendants were conducted by the first author in English, Hindi, Kannada, or Tamil, as the researcher is fluent in these languages. (Note: Participants who selected Telugu as their preferred language during onboarding were interviewed in one of the other four languages.) Interviews on the day of surgery took place in person at the hospital, while post-surgery interviews were conducted either in person at the hospital or through pre-scheduled telephone calls. This decision was influenced by the substantial waiting time on the surgery day—patients typically spend about five hours in the hospital, with the procedure itself taking $\sim$45 minutes. Conversely, during the post-surgery visit, patients receive priority and experience minimal waiting time. Each interview typically spanned 20 to 40 minutes and was audio-recorded with the participant’s consent.
5.2. Procedure for Doctors and Coordinators
We recruited four doctors and two patient coordinators. Three doctors served as operating doctors, and one as both an escalation doctor and knowledge base expert. Both patient coordinators specialize in handling cataract surgery patients, with one serving as a coordinator and the other as an escalation coordinator. The escalation doctor and escalation coordinator held senior positions in the hospital and carried administrative responsibilities. All participating experts attended one of two one-hour training sessions conducted by three researchers. During these sessions, CataractBot was demonstrated, and experts were onboarded. The researchers assumed patient roles, asking medical and logistical questions for the doctors and coordinators to verify and edit. The first session revealed a few bugs, and feedback was collected to enhance the bot. For example, a doctor suggested displaying the question first (rather than the patient’s demographics) in verification messages, as only the top portion is visible in ‘tagged’ WhatsApp messages (Figure 1E). After using CataractBot for two weeks or longer, the participating experts were interviewed by the first author. These interviews followed a semi-structured format, focusing on the bot’s usability, its integration into their workflow, and its impact on interactions with patients. The interviews were audio-recorded, conducted in English, in-person, and typically lasted 45-60 minutes.
5.3. Participants
[…]
5.4. Data collection and analysis
We performed a mixed-method analysis to systematically analyze the collected data comprising of usage logs, interview transcripts, and usability survey responses. CataractBot interaction logs and usability survey data were quantitatively analyzed. Descriptive statistics and statistical tests (such as t-tests and ANOVAs) were used to analyze the count and type of questions asked, verification styles and response times of doctors and coordinators, edits performed by the knowledge base update expert, and responses to the usability survey. Note: For ANOVAs, the sphericity assumption was tested using Mauchly’s test, and in case of sphericity violations, the Greenhouse-Geisser correction was applied. The first author translated and transcribed all interviews (totalling 11.5 hours) into English soon after they were conducted. Throughout the study, all authors regularly engaged in discussions to review observations and interviews. For interview transcripts, we conducted an inductive thematic analysis (Braun and Clarke, 2006) approach, with the first author open-coding all interviews line-by-line. Subsequently, two authors collaboratively reviewed these codes to identify interesting themes in the data (similar to (Okolo et al., 2021)). We iterated on these themes to distill higher-level themes (like “Usage” and “Features”), that we present in our findings below. Note: While there were a total of 49 patients and attendants using CataractBot, at times, we treat them as 31 participants (patient-attendant pairs). This choice stems from instances highlighted during interviews, where patients instructed attendants to query the bot using either of their phones. Additionally, both were interviewed together. Only when comparing their distinct roles, we analyze their usage log data separately.
5.5. Positionality
All authors are of Indian origin, and currently reside in Bangalore, India. One author is a practising ophthalmologist, performing over 10 cataract surgeries weekly, and another author is part of the Quality Control and Patient Safety team at the same hospital. Both were involved in designing the study, shaping research questions, providing feedback on initial bot versions, curating the knowledge base, and ensuring that any risk of harm to the patients was mitigated. The ophthalmologist also served as the knowledge base expert. Two authors specialize in HCI and design, three possess healthcare research experience, and four have expertise in software development. We approached this research drawing on our individual experiences and learning from working at the intersection of computing and healthcare in India. Our interest in making healthcare more accessible to the masses has informed the design of CataractBot and guided our study design.
6. Findings
[…]
Below, Section 6.1 presents the reasons behind patients’ and attendants’ engagement with the bot (answering RQ1). Section 6.2 examines the impact of key features of the bot, including the LLM, experts-in-the-loop and support for multiple stakeholders, languages and modalities (RQ2). Finally, Section 6.3 describes the integration of CataractBot into the doctor-patient workflow and discusses issues of personalization and accountability (RQ3). Table 5 in Appendix A.4 presents a summary of key findings across sections.
6.1. Bot’s Role in Addressing Information Needs
6.1.1. Reasons for (Not) Using CataractBot
Information seekers used CataractBot to address questions they had forgotten or were uncomfortable to ask the doctor or patient coordinator, clarifying answers, and seeking updates. D1 stated that most information seekers come prepared for face-to-face conversations, optimizing their limited time and avoiding multiple visits. CataractBot can alleviate the burden of having to remember every question. Ten information seekers highlighted that the bot helped recall information provided during in-person consultations, which they had either forgotten or not completely understood. Patients found comfort in “written” information, available as a reference (similar to (Jain et al., 2018)), “I need not remember everything.” (A20). Six participants also mentioned using the bot to “double-check” (A7) and seek “reassurance” (A17), despite already knowing certain answers.
Due to existing power differences between doctors and patients (Joseph-Williams et al., 2014) and the judgmental attitudes of some doctors, individuals often hesitate to ask questions. Experts acknowledged that educated individuals are comfortable posing questions. In contrast, “less educated patients who don’t speak English” are the ones for whom CataractBotcan do wonders” (D3). Five participants, all with a Grade-12 education or below, relied heavily on CataractBot asking 57 questions in total, as they reported being unaccustomed to seeking information online. One of them expressed that checking with the bot was more comfortable than approaching a doctor, citing the bot’s non-judgmental nature: “(The patient) has cholesterol issues and heart surgery… the doctor would make faces when I ask about food intake… but not the bot.” (A4).
Despite these reasons, 8 out of 49 information seekers used the bot minimally, sending only one or two messages. Two key reasons emerged. First, six participants mentioned that their in-person interactions with the doctor and patient coordinator sufficed to address their queries. This suggests that CataractBot could play a more pivotal role in primary healthcare centers, where the human infrastructure may be less capable of addressing information needs or where power imbalances make individuals less likely to seek answers. Second, two participants considered cataract surgery to be a “simple” procedure, leading them to have no questions. This indicates that a similar bot could be valuable for more complex (surgical) procedures.
6.1.2. I don’t know” Responses.
Participants complained about receiving “I don’t know” responses (57 total) A total of 36 (10.5%) messages requested status updates—seeking information about the surgery time and discharge time. Due to privacy concerns, CataractBot was not integrated with the hospital’s patient management system, resulting in “I don’t know” responses. Patient coordinators later provided answers. While participants acknowledged the appropriateness of “I don’t know” responses, they suggested providing a “tentative response” (A10) if possible. For instance, to the question “Mine is general anesthesia or local anesthesia?” A10 expected an answer such as “99% of cataract surgeries at Sankara Eye Hospital are done under local anesthesia”. From an expert’s perspective, responding to certain “I don’t know” questions was challenging as they lacked the full context. E.g., A21 asked: ‘I want to postpone the surgery to Feb… is this okay or do I need to get this operated immediately?”, receiving an “I don’t know” response. To correct this, the doctor said: “Please come and check”. This correction frustrated the attendant, who stated: “But (the patient) already went through the complete eye check; the doctor know her condition… I don’t want to come and visit the doctor [again]… it will waste 2-3 hours.” (A21).
When participants received an “I don’t know” response, they sought information from other sources, particularly the internet. For time-sensitive questions, participants reported using the internet to “double-check” the initial LLM-generated, unverified response due to delays in receiving an expert verified response. Four participants, accustomed to relying on the internet for their informational needs, reserved CataractBot exclusively for questions requiring expert opinions, such as “When can i do exercises as anulom vilom?” (P27). According to D4, without CataractBot, patients might have resorted to the internet and risked following inaccurate advice, given the unavailability of doctors for immediate consultations. CataractBot improves the likelihood of receiving accurate responses, as it uses a custom knowledge base.
Figure 3. A. (Left) Messages sent per day pre- and post-surgery. B. (Top right) Messages sent before surgery, on the day of surgery, and after surgery. C. (Bottom right) Message input modality.:
6.2. Effect of Key Features on the Bot’s Usage
6.2.1. LLM-powered
Patients and attendants found the LLM-generated responses from the curated knowledge base were useful, and experts described them as “highly accurate”, noting that they minimized their workload for corrections. These instant responses saved time as information seekers did not have to constantly wait, either physically or over a call. Seven participants emphasized that traveling is time-consuming, and contacting busy hospitals over the phone is challenging due to high patient loads, as “the phone is always busy” (A4). Further, the robustness of the GPT-4 model enabled CataractBot to appropriately handle complex queries from participants with limited literacy, who struggled to formulate questions, consistent with prior research (Jain et al., 2018). For example, P18 asked “Tomorrow operation When should come No phone call received yet?”. CataractBot was able to address 45/65 such questions, of which 26 were asked using audio in an Indic language. Typing in Indic languages using the Latin script (e.g., instead of Devanagari script for Hindi) was rare, occurring in only three messages. In two instances, the messages mixed Hindi and English, and the bot responded accurately:
A8: agar opration k baad pain ho raha to kya karna hai?
Bot: If the patient is experiencing pain post-surgery, they should report to the doctor immediately…
For experts, the LLM enabled them to swiftly correct the bot’s responses using natural language, without worrying about grammatical or spelling errors. For example, in response to a query, D2 added the correction: “Btr to book an appt”. In other instances, experts also instructed the LLM to remove specific sentences, relying on it to generate appropriate responses from their feedback.
6.2.2. Experts-in-the-loop
Information seekers appreciated the novel feature of having human experts verify and update the auto-generated responses (Figure 1F), serving as the key reason for their trust in and engagement with the bot. For instance, “That’s why I gave my WhatsApp (number to onboard the bot), so I can talk directly to the doctor.” (A2). A12 added: “The trust only came when I saw the green tick mark… Before that, I was also in the question mark zone… thinking it’s a machine… But then somebody is checking it, confirming it… It’s actually trustworthy then. Despite our concerns that the icons might be complex, participants found them self-explanatory:
“It’s simple… Even my dad understood. I didn’t tell him [about] the tick or question mark… He told me, “Doctor verified… the tick mark came”… If MY dad was able to use it, I think anyone can.” – A13.
Although the experts-in-the-loop feature increased trust in the final responses, it had a negative impact on perceived bot intelligence. Four participants noted that when an expert corrected CataractBot’s initial answer, their trust in the bot’s unverified answers reduced. For instance, “After some time, when the doctor said it is invalid, I was like, okay, should I even trust the bot?” (P19). In total, 12.8% (44) of the bot’s generated answers (excluding “I don’t know”) for medical and logistical queries were marked incorrect by the expert. This could be due to the limited custom knowledge base, the experts’ high standards for tone and structure of answers, or our prompting strategy favoring caution. Also, since participants knew that these messages were being reviewed by experts, four used the bot to send direct messages to them instead of asking questions (e.g., “Dr. Give me a call?” (P19)).
[…]
Finally, experts did not have access to the final answer sent to information seekers after their suggested edits. This limited verification to a single round, and was designed to minimize their workload. However, this lack of transparency sometimes left experts uncertain if CataractBot had understood their corrections. Only the knowledge base update expert had access to the updated answer sent to the information seeker.
6.2.3. Multimodality
The 19 patients sent 89 messages (43 text, 26 audio, and 20 taps), while the 30 attendants sent 254 messages (163 text, 17 audio, and 74 taps) (Figure 3C). One-way ANOVAs on message types found statistical significance for attendants (F(2, 87)=13.74, p¡.001), but not for patients (F(2, 54)=0.77, p=0.46). Post-hoc pairwise comparisons, corrected with Holm’s sequential Bonferroni procedure, indicated statistically significant differences between ‘text’ and ‘audio’ (t(29)=4.8, p¡.01) and ‘text’ and ‘taps’ (t(29)=3.0, p¡0.01) for attendants. The minimal usage of audio by attendants could be due to demographic differences, with the majority being under 45 years and well-educated. Audio messages were predominantly used by older semi-literate patients. For instance, “I can’t read (or) write Hindi that well, so asking and listening to audio, I did that.” (P12).
Post-surgery limitations on screen use also indicated that speech input and audio output the preferred modality for patients, with 18 out of 26 audio questions by patients asked post-surgery. 31 participants used the related questions at least once. As P7 stated: “I used the audio message once. After that there was no need for me [to type or use audio]… I just kept selecting from the [related questions] options.” While they found the related questions feature convenient and frequently used it, there were instances of dissatisfaction when (in eight cases) the answers to those resulted in “I don’t know” responses. This issue arose because the CataractBot system uses an LLM call to suggest three questions based on the users’ last question, independent of the knowledge base. This approach aimed to uncover gaps in the knowledge base by not restricting questions to the curated knowledge base.
On the expert end, they could only interact with CataractBot using text. This design choice facilitated a “hygiene check” (D4) for thoroughness and precision in experts’ edits, aligning with previous findings (Li et al., 2023).
Patients mentioned utilizing the persistent information provided by CataractBot as a reference and sharing it with others undergoing cataract surgery. Also, as per D1, this information can help when discussing their surgery with experts outside of the Sankara Eye Hospital ecosystem. However, it also increased the sense of accountability among experts, making them “somewhat nervous”. They felt the need to be “very lawyer-like in our conversation because we don’t want it to come back and bite us tomorrow.” (D4).
6.2.4. Multilinguality
The majority of information seekers (29 out of 49) selected English, while 7 opted for Hindi, 6 Kannada, 5 Tamil, and 2 Telugu. None of the participants changed their initially selected language. Given the challenges of typing in Indic languages on smartphones (Joshi et al., 2011), our non-English participants predominantly relied on speech input and selecting from related questions. In total, patients who chose English typed 35 questions, asked 0 questions in speech, and tapped 6 related questions. In contrast, patients opting for Indic languages typed 8 questions, sent 26 audio questions and tapped 14 related questions in total. A similar trend was observed among attendants. Patients found value in local language interaction as input, helping them use the bot independently: “Without even asking me for help, he (patient) was able to use that bot as it supports Tamil.” (A13).
The audio queries in Indic languages were transcribed and translated to English using Azure AI services. However, this was not highly accurate due to the variability in Indic language dialect. This is a known issue and prior work (Jain et al., 2018) relied on human wizards to correct such errors. Experts faced challenges in “deciphering” such messages since they only received the English transcription. In the future, there should be a way to share the original audio query with the expert on demand.
6.2.5. Support for Multiple Stakeholders
[…]
“Dad didn’t try… I was the one thinking about it [the surgery]. He was like, “It’s OK! Not everyone needs to bother in the family… (to know) about what will be this, what will be that”. ” – A3.
Additionally, patients were advised to minimize screen use post-surgery, contributing to reduced bot interaction. Interestingly, patients sometimes offloaded the knowledge-seeking ‘work’ to their active attendants—e.g., “I have been using it. I told him [the patient] what answers I got. So maybe he did not feel the need (to use it).” (A20).
Moving on to experts, both the doctors and patient coordinators highlighted the significant role of escalation experts. Among the 309 verified medical and logistical answers, 62.5% were verified by the operating doctor/coordinator and the rest 37.5% by the escalation expert. Doctors mentioned having 2-3 days per week reserved for surgeries, and on these “operating days”, they rarely checked their phones during work hours, leading to more escalations. Despite their crucial role, escalation experts interacted differently from operating doctors and coordinators. Unlike the latter, who had face-to-face meetings with patients, escalation experts lacked contextual knowledge, limiting them to providing generic responses.
6.2.6. Knowledge Base
Out of the 91 responses edited by experts, 72 (79.1%) received approval from the knowledge base expert (D4) to be added to the knowledge base. For 49.3% of these, the bot’s final response (generated by merging the initial response and expert edit suggestion) underwent additional editing by the knowledge base expert. These changes were minimal with an average relative Hamming distance of 28.2%, and mainly addressed aspects of “tone and structure”, such as preferring active voice and making responses more generally applicable across patients. This hints that the bot effectively incorporated experts’ edits. Between the first and sixth weeks of the study, the proportion of LLM-generated answers that experts marked as “accurate and complete” increased by 9.9% for medical questions and 22.9% for logistical questions. This suggests that as the knowledge base was updated, the experts’ verification workload decreased.
6.3. Bot Integration into Doctor-Patient Workflows
6.3.1. Workflows
The majority of messages (71.4%) were sent in the 8-day period from 5 days prior to surgery to 2 days post-surgery with the highest number of messages (76) sent on the surgery day (Figure 3B). Messages were sent throughout the day (7am to 11pm), as A18 mentioned: “Whenever I am idle at work or at home… whatever (question) comes to my mind, I just ask. The workflow employed by information seekers varied. Most participants would ask a question, leave their phones, and check the answer later when free. They hardly noticed the wait due to being occupied with their day-to-day activities. In such cases, notifications played a key role. The message stating that the expert had provided a verified response (Figure 1G), made the otherwise unnoticeable green tick (Figure 1F) prominent. Participants, when in need of a quick response, either relied on the initial (LLM-generated) answer or re-verified it on the internet.
[…]
“I know that it will take its own time… based on their [doctors’] busy schedule, their availability… It’s not like there’s some dedicated doctor monitoring it all the time, right? If that is the case, then (the) bot is not required. They [doctor] can directly reply.” – A17.
Notifications played a key role. The message stating that the expert had provided a verified response (Figure 1G), made the otherwise unnoticeable green tick (Figure 1F) prominent and the bot’s process self-explanatory.
For experts, the asynchronous nature of the bot provided them the flexibility to offer verifications at their convenience. The majority of responses (31) were provided by experts at 4-5 pm, post their busy workday, as most OPD and surgeries finish by 4 pm. The second-highest responses (28) occurred at 1-2 pm, during their lunch break. This usage pattern was described as: “I use my mobile phone only when I go for breaks… lunch break, coffee break, or in the morning before I start OPD. The rest of the time, my (cellular) data is off.” (D1). Experts refrained from using their phones in front of patients due to the negative perception associated with using phones for personal reasons. We found that the experts were “fine” with the additional task of verifying the bot’s responses, as articulated by D4: “It takes only about 15-20 seconds to answer a question… It was not really eating into my time… It’s much easier than replying to my mailbox.
As five experts used CataractBot on their personal WhatsApp accounts, it blurred the boundary between their home and work, similar to previous findings (Mols and Pridmore, 2021; Wang et al., 2020). Our experts also noted that WhatsApp had a low signal-to-noise ratio, resulting in instances where they missed CataractBot messages: “My WhatsApp has 5000+ [unread] messages… college groups, family groups… I just see messages in the important groups.” (D2). Addressing this issue, the escalation doctor D4 suggested that experts should be recommended to “pinCataractBot on WhatsApp, ensuring it remains consistently at the top of their chat list.
6.3.2. Personalization vs Privacy
Three participants valued that the bot could not access their medical records.
“Using personal health information [by the bot] is… a violation of patients’ rights. If their data is added… external agencies can extract that information and might use it for advertising.” – A17.
These participants were educated (holding Master’s degrees) and considered personalization as “an unnecessary add-on”. They raised concerns about complications associated with sharing PII with LLMs, both in terms of legislation and “privacy threats”. A1, for example, remarked “I would definitely not want my mom’s data to be out there, all over the internet.”. Moreover, lack of personalization enabled usage of the chatbot for others undergoing cataract surgery: “Two people I know are going to have their cataract surgery soon… I’ll use this bot for them.” (A11).
On the other hand, fueled by the recent hype around AI and LLMs, eight participants expected “personalized” answers to their queries with little concern for privacy, and were slightly disappointed with generic responses. This aligns with prior findings indicating minimal privacy concerns related to digital data in the global south, specifically in India (Okolo et al., 2021). Interestingly, three participants did not even consider their medical data to be “private”, as they stated, “It’s nothing personal to be hidden. All (of us) have common health issues.” (A14). Moreover, two participants questioned why others would be interested in their medical data, as “it is not useful for others” (P18).
Two participants wanted personalization, along with the sharing of patient data, to be an “opt-in” feature where individuals conduct a cost-benefit analysis and make an informed choice based on their preferences. They believed that a personalized bot would offer more relevant, and fewer “I don’t know”, responses.
“If you don’t want to share your information, it (the bot) should say “If you ask personal questions, I won’t have access, so you may have to visit the hospital.”… We have to compromise.” – A21.
Five experts voted to integrate the chatbot with patient data, including schedules and consultation history, to offer personalized responses. D3 suggested integrating patient medical history to facilitate customized responses from the bot and doctors: “We don’t even know whether they’re diabetic, they’re hypertensive. So, when they ask “What should I eat and come?”… I am not sure what to say.” (D3). However, D1 and D2 also voiced skepticism towards the LLMs capability to understand complex patient records accurately. “There are a lot of variables… the surgeon, the patient, their tests… It’s too much information to process.” (D1). In addition, the knowledge base expert cautioned that highly personalized responses would result in minimal updates to the knowledge base, potentially increasing CataractBot-related workload for experts and negatively impacting chatbot adoption among experts.
7. Discussion
In this paper, we present CataractBot, an LLM-powered experts-in-the-loop WhatsApp chatbot to address the information needs of patients undergoing cataract surgery. Building on limited prior research in multi-stakeholder settings (Yang et al., 2024; Jo et al., 2023), we provide valuable insights into LLM-mediated interactions among patients, attendants, doctors, and coordinators. Our field study revealed positive evidence of CataractBot’s usefulness and usability across stakeholders. Doctor-verified responses were key to patients’ and attendants’ trust in and engagement with the bot. They appreciated that CataractBot instantly answered questions anytime, saving time by reducing the need for hospital calls or visits. Those with limited (tech-)literacy found it easy to use because of the multilingual and multimodal support. Doctors and coordinators commended CataractBot for acting as a facilitator, creating a layer of privacy between them and patients, and providing them the flexibility to verify responses at their convenience.
Below, we identify design implications, raising critical insights and open questions that warrant further attention for such experts-in-the-loop chatbots to succeed on a larger scale across domains.
7.1. Design Implications
7.1.1. Generalization and Scaling
We designed and evaluated our experts-in-the-loop chatbot system in the context of cataract surgery in India. Expanding it beyond India may require further research. In the United States, for example, insurance plays a crucial role in healthcare. Patients often require prior authorization for treatments, and they would value a bot providing information in navigating these processes (Sun et al., 2013). This necessitates involving insurance-specific experts in knowledge base creation and response verification. Also, our study raised concerns about expert accountability due to the persistence of written communication in the bot. This issue could be more pronounced in the Global North, where doctors are increasingly wary of lawsuit risks (Sun et al., 2013). Balancing these needs and tensions will be critical for CataractBot’s adoption across geographies. Finally, health communication is most effective when users perceive the source as demographically and attitudinally similar to themselves, not just contextually relevant (Kreuter and McClure, 2004). This highlights the need to adapt CataractBot for different settings by (1) parametrizing context-specific details, such as payment methods and hospital protocols, and (2) tailoring language to ensure cultural relevance through familiar terms and metaphors.
Our open-sourced framework can be adapted beyond cataract surgery, to support more complex, atypical, or long-term treatments–e.g., glaucoma surgery, cancer treatment, or pregnancy–where easy access to verified information would be valuable for patients. The applicability can be extended further, to fields beyond the medical domain, such as law, finance and education. In these scenarios, end-users (e.g., defendants, taxpayers, or students) could receive synchronous responses from an LLM using a custom knowledge base, and subsequent asynchronous verified responses from experts (e.g., lawyers, accountants, or teachers). Expert-mediated bots can even be used to upskill less-trained workforce members. For instance, these bots could serve as valuable learning resources for community health workers, who typically have minimal medical training (Yadav et al., 2019).
7.1.2. Expert-in-the-loop Approaches
Prior studies have raised concerns about the application of LLMs in healthcare scenarios, due to their inconsistency, potential errors, and bias (Denecke et al., 2024; Au Yeung et al., 2023). Moreover, users who are less (technologically) literate face a heightened risk of harm (Kapania et al., 2022). Our experts-in-the-loop framework not only addresses these concerns but also adheres to OpenAI’s usage policies (OpenAI, 2024), which state that “tailored medical/health advice cannot be provided without review by a qualified professional”. By using our framework, chatbots can operate at the “sweet spot of patient-LLM-clinician collaboration” (Hao et al., 2024), where the capabilities of technology and human judgment are integrated. In our study, this ensured reliability and cultivated trust, as information seekers knew that their operating doctor was verifying responses. However, trust in the initial automated answers reduced when these answers were subsequently marked as incorrect/incomplete by the doctor. To address this in future work, one design approach could be to eliminate the initial automated response, providing only expert-verified answers. We caution that this may lead to long wait times, potentially discouraging use. For high-risk use cases like cancer treatment,verified accuracy may be prioritized over immediacy. Future research could compare the two paradigms: a hybrid system with synchronous automated answers followed by asynchronous verification versus a fully asynchronous system delivering only verified answers. Such a study would provide valuable insights into balancing immediate access with the accuracy required for critical medical information in chatbot design.
Doctors are a scare resource in healthcare ecosystems. Our chatbot system can help scale their impact, by reducing the time spent answering routine questions at length during consultations and calls, thus freeing up time for additional patient care. However, as doctors had to verify responses during deployment, CataractBot also added to their workload. To minimize this, one approach is to deliver pre-verified answers to specific questions that the expert has repeatedly marked as correct. Given the potential for LLM hallucinations (Denecke et al., 2024), cached responses must be carefully matched to similar patient demographics and question phrasing before reuse. Another approach is to use a second LLM for quality control while generating each response, as GPT-4’s evaluations of AI-generated responses align well with human expert assessments (Gumma et al., 2024). This maker-checker verification module would enable tracking of the bot’s response quality, especially as the knowledge base expands. The module could also evaluate the confidence level in the generated response, or the perceived urgency of the question, which could be used to prioritize sending nudges to experts for verification.
7.1.3. Multimodality and Personalization
Our system faced challenges in accurately interpreting questions from voice messages due to limitations in transcription and translation technologies. The same inaccurate transcriptions were shared with doctors leading to confusion. Such issues would not arise in deploying a voice-based input modality to a predominantly English-speaking population or a tier-1 language setting in the Global North, as seen in prior work (Yang et al., 2024). Additionally, non-verbal cues in audio messages–such as signs of anxiety, impatience, or the presence of others–were lost in textual transcriptions. This hindered experts’ ability to deliver appropriate care (also noted in (Li et al., 2023)). To address this, we suggest making the original audio clips available to experts on-demand. Beyond audio, considering the visual capabilities of recent LLM models like GPT-4V, such bots could be enhanced to process image-based inputs, aligning with recommendations in prior studies (Jain et al., 2018). For example, a patient could upload a photo of a specific medicine to ask questions about it, instead of providing a written description. While this could improve the user experience for information seekers, it may increase the workload for experts. Further, prior research (Hao et al., 2024) cautions that although LLMs are proficient in text processing, their ability to handle other forms of media, such as image and video, remains less reliable.
Moving on to personalization, while tailored responses can boost adoption and utility by providing more relevant information, they also raise privacy concerns, potentially reducing usage among certain users. This issue may be more pronounced in the developing world, where there is more awareness around privacy. Additionally, experts may face increased workloads as they reference user-specific information for verification, since many personalized edits may not be suitable for inclusion in the generic knowledge base. Low (tech-)literate users might require assistance in making informed decisions about the use of their data (Okolo et al., 2021). A binary approach of either sharing all records or none may not address individual preferences. Instead, users should have the agency to selectively share records they deem relevant and acceptable for the bot’s access. This aligns with the philosophy of India’s National Digital Health Mission (National Health Authority and Family Welfare, 2020). One solution could be to allow users to upload relevant documents or manually enter specific details they want the bot to access. This can empower users to control shared information while benefiting from a personalized experience. Further, when developing CataractBot in May 2023, we chose to use GPT-4, as it was the leading LLM (OpenAI et al., 2023). However, recent open-source models like LLaMA have shown comparable performance to GPT in medical applications (Li et al., 2025). Incorporating these models into the CataractBot system would enable local processing of patient data, helping address privacy concerns.
7.2. Limitations
We acknowledge several limitations of this work. First, results should be interpreted with caution, given the specificities of the study, including a relatively small sample size, users’ first encounter with a medical chatbot, and the cataract treatment scenario. Positive user responses, minimal increase in expert workload, and other findings should be validated in future studies, preferably with longitudinal designs. Second, system-level analyses, such as scalability testing and power measurement, are needed before deploying the proposed solution more broadly. Finally, although participants reported high trust, repeated “I don’t know” responses or inaccurate information during regular use may erode this trust. We note the potential risk of disseminating inaccurate medical information at scale with chatbot and LLM technologies. Therefore, stronger methods are needed to develop, review, and update the knowledge base, as well as to routinely evaluate the accuracy of bot’s responses.
8. Conclusion
Building a system that addresses patients’ information needs by providing expert-verified information is an open problem in healthcare. We propose a novel solution—an LLM-powered experts-in-the-loop chatbot framework, that utilizes retrieval-augmented generation over a custom knowledge base to provide synchronous responses, and utilizes experts to provide verified responses asynchronously. Our in-the-wild study involving 49 information seekers (patients and attendants) and 6 experts (4 doctors and 2 coordinators) demonstrated the positive impact of this technological intervention. Patients not only trusted the chatbot’s response, but were also willing to wait, understanding that their busy doctors were verifying them. Simultaneously, doctors and coordinators, despite their hectic schedules, made time during breaks and at home to help patients by verifying LLM-generated responses. The favorable feedback from various stakeholders in the healthcare ecosystem indicates that a chatbot, delivered through ubiquitous smartphones and WhatsApp, can effectively enhance information access in critical healthcare settings, even for individuals with limited literacy and technology experience.
Acknowledgements.
Many thanks to all the participants for their time and patience.
Appendix A Appendix
A.1. CataractBot Implementation Details
A.1.1. CataractBot LLM Prompts
The CataractBot system leverages LLM (GPT-4 in our case) for these four tasks (Table 2):
  1. (1)
    Response Generation: For every medical and logistical question asked by the patient/attendant, the system performs a vector search on the Knowledge Base (KB) to extract the three most relevant data chunks related to the query. The LLM is then prompted (Table 2) to extract an answer for the query from these data chunks. Note: Question Classification is part of the same LLM call to improve efficiency.
  2. (2)
    Related Questions Generation: For every medical and logistical question, the system prompts LLM to generate three related questions based on the preceding query. Note: The 72 character limit is due to the WhatsApp’s message limit in interactive suggestions.
  3. (3)
    Final Response Generation: If the expert marks the initial LLM-generated response as incorrect or incomplete, the system prompts LLM to generate the final response by merging both the initial response and the expert’s edit suggestion.
  4. (4)
    Shorten Response: If the generated response exceeds WhatsApp’s message limit of 700 characters, the system prompts LLM to summarize it within the specified character limit.
Table 2. LLM Prompts used in CataractBot.
System Prompt Query Prompt
Response
Generation
You are a Cataract chatbot whose primary goal
is to help patients undergoing or undergone a
cataract surgery. If the query can be truthfully
and factually answered using the knowledge
base only, answer it concisely in a polite and
professional way. If not, then just say “I do
not know the answer to your question. If this
needs to be answered by a doctor, please
schedule a consultation.”
In case of a conflict between raw knowledge
base and new knowledge base, prefer the new
knowledge base. One exception to the above
is if the query is a greeting or an acknow-
ledgement or gratitude. If the query is a
greeting, then respond with a greeting. If the
query is an acknowledgement or gratitude to
the bot’s response, then respond with an
acknowledgement of the same. Some exam-
ples of acknowledgement or gratitude to the
bot’s response are “Thank You”, “Got it”
and “I understand”. In addition to the above,
indicate like a 3-class classifier if the query
is “medical”, “logistical” or “small-talk”.
Here, “small-talk” is defined as a query
which is a greeting or an acknowledgement
or gratitude. Answer it in the following
json format:
The following knowledge base have been
provided to you as reference:
Raw documents are as follows:
<relevant chunks string>
New documents are as follows:
<relevant updated chunks string>
The most recent conversations are here:
<conversation string>
You are asked the following query:
<user query>
Ensure that the query type belongs to only the
above mentioned three categories. When not
sure, choose one of “medical” or “logistical”.
Related Questions
Generation
What are three possible follow-up questions
the patient might ask? Respond with the
questions only in a python list of strings. Each
question should not exceed 72 characters.
A patient asked the following query:
<query>
A chatbot answered the following:
<response>
Final Response
Generation
You are a Cataract chatbot whose primary
goal is to help patients undergoing or under-
gone a cataract surgery. A cataract patient
asks a query and a cataract chatbot answers
it. But, the doctor gives a correction to the
chatbot’s response. Update the cataract
chatbot’s response by taking the doctor’s
correction into account. Respond only with
the final updated response.
A cataract patient asked the following query:
<query>
A cataract chatbot answered the following:
<response>
A doctor corrected the response as follows:
<correction>
Shorten
Response
You are a Cataract chatbot, and you have
to summarize the answer provided by a bot.
Please summarise the answer in 700 char-
acters or less. Only return the summarized
answer and nothing else.
You are given the following response:
<response>
A.1.2. CataractBot’s Support for Multimodal Communication
Figure 4. A question asked (using audio), receiving an unverified response and Related Questions from CataractBot.:
See Figure 4.
A.2. Participant Demography
Table 3. Demographic details for 19 patients and 30 attendees, with their interview participation, the duration of CataractBot usage (calculated as the difference between the first and last day of messages sent), and the total number of messages sent.
PId Age Sex Language
Highest
Education
AId Age Sex Language
Highest
Education
Surgery-day
Interview
Post-surgery
Interview
Days
# of
Messages
A1 43 M English Masters Yes Yes 7 9
A2 31 F Kannada Grade 12 Yes Yes 3 13
[…]
P4 52 M English Grade 12 A4 39 F English Grade 12 Yes No 4 8
[…]
A8 39 M Hindi Bachelors Yes No 6 6
P9 69 M English Masters A9 38 F English Masters Yes Yes 6 14
A10 71 M English Bachelors Yes Yes 26 15
[…]
P12 65 M Hindi Bachelors A12 31 F English Masters Yes Yes 2 6
P13 69 M Tamil Grade 12 A13 35 M English Masters Yes Yes 25 18
[…]
A15 57 M English Grade 12 Yes No 4 4
[…]
A17 39 M English Bachelors Yes Yes 2 8
P18 56 M English Grade 12 A18 23 M English Masters Yes Yes 27 24
P19 41 M English Masters A19 36 F English Bachelors Yes No 19 15
P20 64 M Tamil Grade 12 A20 60 F Tamil Grade 12 Yes No 4 16
A21 43 M English Masters Yes No 15 19
A22 35 F English Bachelors Yes No 7 11
P23 54 M English Bachelors 7 8
P24 66 M Kannada Bachelors A24 37 M English Masters No No 1 1
[…]
P26 65 M Hindi Bachelors A26 34 M English Bachelors No No 19 15
P27 64 M Hindi Bachelors A27 35 M English Bachelors No No 18 14
A28 40 M English Masters No No 1 1
[…]
P30 50 M Hindi Masters A30 25 F English Bachelors No No 5 3
P31 63 M Kannada Bachelors A31 32 M English Bachelors No No 15 36
Table 4. Demography details of doctors and patient coordinators who participated in the formative and/or deployment study
Id Role(s) Age Sex
Highest
Education
Experience
Surgeries
per week
Formative
study?
Deployment
study?
D1 Operating doctor 42 F Masters 10+ years 10-20 No Yes
D2 Operating doctor 44 F Masters 15+ years 10-20 No Yes
D3 Operating doctor 46 M Masters 20+ years 30-40 Yes Yes
D4 Escalation doctor, Knowledge base expert 46 M Masters 20+ years 15-20 Yes Yes
D5 Doctor 30 M Masters 6+ years 10 Yes No
D6 Doctor 29 F Masters 2+ years 1-2 Yes No
C1 Operating patient coordinator 33 F Bachelors 10+ years N/A Yes Yes
C2 Escalation patient coordinator 36 F Bachelors 15+ years N/A No Yes
C3 Patient coordinator 45 F Bachelors 10+ years N/A Yes No
See Tables 3 and 4.
A.3. Chatbot Usability Evaluation Form
Responses were given on a 5-point Likert scale, ranging from Strongly Disagree to Strongly Agree.
  1. (1)
    CataractBot understands me well.
  2. (2)
    CataractBot’s responses are easy to understand.
  3. (3)
    CataractBot’s responses were useful, appropriate, and informative.
  4. (4)
    CataractBot responds quickly.
  5. (5)
    CataractBot seems to have a good grasp of medical knowledge.
  6. (6)
    CataractBot is kind and helpful.
  7. (7)
    CataractBot is easy to use.
  8. (8)
    I would be willing to use CataractBot (or a similar bot before/after a major surgical treatment) in future.
A.4. Summary of Key Findings
Table 5. Key findings from CataractBot deployment study
Research Q Theme Finding (Information Seekers) Finding (Experts)
RQ1: Information Needs Reasons for usage
• Asking forgotten or uncomfortable questions
• Clarifying and verifying information
• Seeking updates
• Commitment to patients
Reasons for
lack of usage
• In-person expert interactions suffice
• Lack of questions
• “I don’t know” responses
• Persistent medium gives rise to
accountability concerns
RQ2: Features LLM-powered
• Instant answers.
• Saves time.
• Understands complex/ill formed questions
• Accurate answers
• (Informal) corrections are quick and easy
Experts-in-the-loop
• Drives trust and engagement
• Lowers perceived intelligence of bot
• Adds a layer of privacy from patients
• Corrections expand bot’s answers
• No access to final answer reduces both
workload and transparency
Multimodality and
multilinguality
• Independent usage by older,
less educated, visually impaired patients
• Hard to decipher English transcriptions
Support for
multiple stakeholders
• Attendants (younger, educated) message
more than patients
• Escalation experts are crucial, but
lack patient context
Knowledge base
• Minimal edits after first expert verification
• Bot’s performance improves over time
• Experts’ workload eases
RQ3: Workflow Workflows
• Messages spread throughout hours of day
• Most messages on surgery day
• Waiting for verification is fine
• Verify at their convenience
• WhatsApp blurs work-home boundaries
Personalization
vs. privacy
• No personalization implies limited usefulness
• Medical history not seen as private by some
• Personalization wanted on opt-in basis
• Lack of medical history makes verification
challenging
• Personalization would add complexity
and workload
See Table 5